### v1

# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

Abstract
In this paper, we address a significant gap in the field of neuroimaging by introducing the largest-to-date public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in Volume Electron Microscopy (VEM) images. The intricate relationship between cerebral blood vessels and neural function underscores the vital role of vascular analysis in understanding brain health. While imaging techniques at macro and mesoscales have garnered substantial attention and resources, the microscale VEM imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure. As researchers delve deeper into the microscale intricacies of cerebral vasculature, our BvEM benchmark represents a critical step toward unraveling the mysteries of neurovascular coupling and its impact on brain function and pathology. The BvEM dataset is based on VEM image volumes from three mammal species: adult mouse, macaque, and human. We standardized the resolution, addressed imaging variations, and meticulously annotated blood vessels through semi-automatic, manual, and quality control processes, ensuring high-quality 3D segmentation. Furthermore, we developed a zero-shot cortical blood vessel segmentation method named TriSAM, which leverages the powerful segmentation model SAM for 3D segmentation. To lift SAM from 2D segmentation to 3D volume segmentation, TriSAM employs a multi-seed tracking framework, leveraging the reliability of certain image planes for tracking while using others to identify potential turning points. This approach, consisting of Tri-Plane selection, SAM-based tracking, and recursive redirection, effectively achieves long-term 3D blood vessel segmentation without model training or fine-tuning. Experimental results show that TriSAM achieved superior performances on the BvEM benchmark across three species.
1 Introduction
With around 2% of body weight, our brain receives around 20% of blood supply. With inadequate blood supply, such as stroke, an average patient loses $\sim$1.9 million neurons (Saver, 2006). Alterations of blood vessel structures are observed in many brain diseases, such as Alzheimer’s and vascular dementia (Kalaria, 2010). Thus, blood vessels in the brain have been extensively investigated by a variety of imaging methods with different resolutions and structural details (Figure 1a). Compared to the macro-level imaging (e.g., CT and MRI (Dyer et al., 2017; McDonald & Choyke, 2003)) and mesoscale-level imaging (e.g.light microscopy Lochhead et al. (2023)), Volume Electron Microscopy (VEM) (Peddie & Collinson, 2014; Peddie et al., 2022) can reveal the detailed ultrastructure including endothelial cells, glial cells, and neurons (Figure 1b).
Traditionally, the imaging methods at the macro and mesoscale are widely used and have produced a large amount of data, and a variety of image segmentation algorithms, public datasets, and evaluation methods have been developed  (Goni et al., 2022; Moccia et al., 2018). Recently, owing to the rapid improvement of imaging technology, the sample size of VEM is significantly increased covering all the layers of the cerebral cortex of mouse (The MICrONS Consortium et al., 2023)and human brain (Shapson-Coe et al., 2021), as well as the whole brain of fly (Dorkenwald et al., 2023). However, to robustly analyze blood vessels in VEM images, existing segmentation methods suffer from two major challenges: the diversity of the image appearance due to variations in the involved imaging pipeline including sample preparation, and the complexity of blood vessel morphology.
Figure 1: Comparison of imaging modalities for blood vessel analysis. (a) Both microtomography ($\mu$CT) Dyer et al. (2017) and light microscopy (LM) Lochhead et al. (2023) can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details. (b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all the cell types and their details are visible for further investigations.
To accelerate the method development, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images. The VEM image volumes of the BvEM dataset are from recent publications, which are the largest for each of the three mammal species: mouse, macaque, and human acquired at different VEM facilities. We downsampled the volumes to a consistent resolution and performed extensive blood vessel annotation, including manual proofreading, semi-automatic segmentation error correction, and quality control, involving multiple rounds of scrutiny by neuroscience experts to ensure accuracy and completeness.
We then propose a zero-shot 3D segmentation method named TriSAM based on the Segment Anything Model (SAM) (Kirillov et al., 2023) which is able to segment objects in the image given points or a bounding box as the input prompt. With a multi-seed tracking framework, our proposed TriSAM consists of three components: Tri-Plane selection, SAM-based tracking, and recursive redirection. Given the fact that tracking along blood vessel flow direction is easier, our Tri-Plane selection method chooses the best plane for tacking. To further exploit the 3D blood vessel structure and ensure long-term tracking, we propose to change the tracking direction at the potential turning points with a recursive redirection strategy. The proposed method achieves state-of-the-art performance compared to the baseline methods on the proposed BvEM benchmark across all three species.
2 BvEM Dataset
Dataset Design. To foster new methods tackling the challenges, we built our BvEM dataset on top of the largest publicly available VEM image volumes for samples from each of three mammal species: visual cortex from an adult mouse The MICrONS Consortium et al. (2023), superior temporal gyrus from an adult macaque Loomba et al. (2022), and temporal lobe from an adult human Shapson-Coe et al. (2021). Each dataset was acquired with different protocols at different VEM facilities and we refer readers to the respective papers for more details.
Image Volume Standardization. We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and a manageable dataset size for computation and dissemination for AI researchers. Then, we trimmed off the rows and columns across all slices near the image boundary of the mouse and human data, where the blood vessels are hard to annotate due to the missing image content. We directly used the macaque dataset which was well-trimmed by its authors. As shown in Figure 2, the imaging quality and the appearance of blood vessels vary drastically across these three datasets, showcasing the cutting-edge large-scale VEM pipelines in the field.
Figure 2: BvEM Dataset. We compiled the largest publicly available VEM dataset for each of the (a) mouse The MICrONS Consortium et al. (2023), macaque Loomba et al. (2022), and human Shapson-Coe et al. (2021) samples acquired by three different VEM labs. For each dataset, we downsampled the image volume to a similar resolution and trimmed off boundary regions. Then, we corrected the existing ground truth or instance segmentation results for each blood vessel instance (displayed in different colors).
[…]
Table 1: Dataset Statistics. Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined.
Sample Microscope Resolution (nm) Volume size (voxel) Length: max/sum (mm)
Mouse TEM 320$\times$ 256$\times$256 2495$\times$3571$\times$2495 1.6/1.7
Macaque SBEM 240$\times$176$\times$176 450$\times$1271$\times$995 713.3/714.5
Human MultiSEM 264$\times$256$\times$256 661$\times$7752$\times$13500 107.2/126.7
Dataset Statistics. We summarize the dataset details in Table 1, where the dimension is in the “zyx" order. The BvEM-Macaque data has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human image data are around 80 and 121 times bigger than it. From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines with the Kimimaro software (Silversmith et al., 2021) and computed the length for each blood vessel instance. Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each dataset. Note that the BvEM-Human dataset is a thin slab, where many blood vessels are disconnected due to the limited spread.
3 Method
Conventional blood vessel segmentation heavily depends on a substantial volume of manually annotated data, a resource that is notably scarce in the existing literature. Therefore, we propose a zero-shot 3D blood vessel segmentation method based on SAM (Kirillov et al., 2023) which achieves superior performance without model training or fine-tuning.
Figure 3: Problem formulation and challenges. (a) The 3D segmentation task can be formulated as the multi-seed object tracking problem. There are two main challenges for such an approach: (b) generating high-quality initial seeds (blue dots) and minimizing the false positive (red crosses) and false negative errors (yellow circles); (c) adaptively selecting tracking directions as the quality of SAM segmentation results can vary significantly.
3.1 Problem formulation: 3D Segmentation as multi-seed tracking
To apply SAM to 3D VEM images, a naive method is to combine SAM with a tracking algorithm. However, blood vessels are sensitive to split error requiring robust long-term tracking which is a remaining challenge in object tracking literature. Another possible solution is to use a 3D adaptor with 2D SAM features which requires a large amount of labeled data for training. However, labeled training data is limited since the annotation for VEM images needs the involvement of experts. To this end, we formulate 3D segmentation as a multi-seed tracking framework shown in Figure 3a that doesn’t require any annotated data or robust long-term tracking algorithms. To achieve long-term tracking with multi-seed tracking, we leverage the fact that one of the planes is reliable for tracking while other planes indicate potential turning points. In particular, different from videos, the VEM data can be interpolated from different planes (e.g. yx/zy/zx planes). Tracking along some planes is extremely difficult since the shape and scale change dramatically. Therefore, it is essential to select which plane should be used for tracking. To this end, we propose a Tri-Plane Selection method to select the best plane for tracking based on the potential blood flow direction. To achieve long-term tracking of blood vessels, a Recursive Redirection method is proposed to keep tracking at potential turning points. Based on these two components, a simple SAM-based Tracking is effective in achieving long-term 3D blood vessel segmentation.
3.2 TriSAM 3D Segmentation
The proposed TriSAM 3D segmentation method has three components: Tri-Plane selection, SAM-based tracking, and recursive redirection. The initial seeds are first generated automatically or manually. Then, TriSAM 3D segmentation is performed on each seed and the segmentation results are combined as the final segmentation. The pipeline of the proposed method is shown in Figure 4.
Initial Seed Generation. Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background. Therefore, pixels darker than a threshold are considered as background while others are seed candidates. To improve efficiency, we only keep the center of each connected component as the final seeds. In practice, the threshold is set to $\eta I_{m}$ where $I_{m}$ is the maximum intensity of the volume.
Tri-Plane Selection. Rather than videos, 3D VEM images can be interpreted in different planes. In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically. As shown in Figure 4, tracking along the bottom plane is hard because the shape is complex and changes quickly. This phenomenon can be further confirmed by the experiment shown in Figure 7 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis. Tracking along the plane according to the blood flow direction is ideal and this can be achieved based on choosing a plane with smaller segmentation as shown in Figure 4. Furthermore, we take into account the probability associated with the SAM segmentation result, as it reflects the confidence level in identifying blood vessels. Overall, the plane with the smallest segments and a probability higher than the threshold $\tau$ is selected as the best plane.
SAM-based Tracking. By tracking along the optimally selected plane, changes in shape and size of the blood vessel are more stable compared to other planes. Therefore, we use the enlarged bounding box and the center of the segment as the prompts for the next slice. In particular, we first pad the bounding box of the segment as the bounding box prompt and compute the center of the segment as the point prompt. These prompts are used for segmenting blood vessels in the next slice. The tracking will stop when the probability of the segment is less than a threshold $\tau$.
Figure 4: Left: SAM-based tracking. Right: our proposed TriSAM. (a) Tri-Plane selection is first proposed to select the best plane for tracking. (b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking direction. (c) Recursive redirection exploits potential turning points for long-term tracking.
Recursive Redirection. SAM-based tracking is simple and fast for short-term tracking. But it may fail when shape and scale change dramatically as shown in Figure 3c. To enable long-term tracking, we propose a recursive redirection method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels. Specifically, for each seed, we maintain a turning point tree where the root is the initial seed. In each iteration, we sample a left node in the turning point tree and perform Tri-Plane selection and track along the chosen plane. As blood vessels may change directions within this iteration, we then split this node to obtain new potential turning points. We find that the non-best planes potentially identify turning points or endpoints accurately as shown in Figure 4, Therefore, we propose to generate the turning points according to the segmentation results at non-best planes. In particular, when tracking along the z-axis, we apply SAM to segment blood vessels in the zx and zy planes. Potential turning points are identified as points in these segments with the smallest and largest z-values.
4 Experiments
In this section, we present TriSAM’s experimental results. We begin with settings and implementation details. Then we show the benchmark results. Finally, we perform ablations for further analysis.
4.1 Experimental Settings and Implementation Details
Dataset Split. We used the whole dataset for evaluation. For the supervised setting, we randomly sampled subvolumes are annotated. We sampled 4 subvolumes of size 256$\times$256$\times$256 for the BvEM-Macaque dataset amounting to around 10% of the dataset, and 12 subvolumes of size 384$\times$384$\times$384 for each of the BvEM-Mouse and BvEM-Human datasets amounting to around 1% of each dataset. These annotated subvolumes were divided into a 3-1 train-valid split.
Implementation Details. Since the volume is too large, non-overlapping subvolumes are used for segmentation and the results are fused to form the final prediction. For BvEM-Macaque, we directly predict the whole volume. For others, $k\times 1024\times 1024$ subvolumes are used since $1024\times 1024$ is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse. For preprocessing, we utilize defliker method and temporal smoothing on the z-axis to remove the artifacts and missing slices. For postprocessing, holes are filled and small connected components are removed after the segmentation. Global color thresholding is used to generate the initial seeds and $\eta$ is set to 0.98 as shown Figure 7 (left). Given the significantly greater complexity of human blood vessels, we employ the manually generated seeds for the challenging BvEM-Human dataset. For SAM-based tracking, the probability threshold $\tau$ is set to 0.8 according to the experiment shown in Figure 7 (right). Unless explicitly stated otherwise, we use MobileSAM (Zhang et al., 2023) instead of the standard SAM (Kirillov et al., 2023) in all conducted experiments to improve the inference speed. All experiments are conducted on an NVIDIA-A100 GPU.
Evaluation Metrics. Following Weigert et al. (2020), we use Precision, Recall, and Accuracy as the metrics to evaluate the performance:
$Precision=\frac{TP}{TP+FP},\quad Recall=\frac{TP}{TP+FN},\quad Accuracy=\frac{TP}{TP+FP+FN},$ (1)
where $TP$, $FP$, and $FN$ are instance-level true positive, false positive, and false negative respectively. We use instance-level metrics since it is more sensitive to split errors. In particular, the Hungarian algorithm is used to match ground-truth instances and predicted instances with negative Accuracy as the cost matrix.
Table 2: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the amount of proofreading effort. We compare the proposed TriSAM with various baseline methods under fully supervised with limited training data, unsupervised, and zero-shot settings.
Method Training Data BvEM-Mouse BvEM-Macaque BvEM-Human
Pre Rec Acc Pre Rec Acc Pre Rec Acc
Initial Annotation  [^1] Results from original dataset papers. BvEM-Mouse and BvEM-Macaque are automatically segmented while BvEM-Human is manually labeled. Unknown 93.74 36.62 35.74 1.65 23.17 1.57 100.00 25.68 25.68
3D UNet Limited 13.45 0.91 0.67 16.04 89.65 15.75 68.63 2.46 2.43
nnUNET Limited 3.54 49.55 6.59 24.34 29.57 23.74 3.44 23.20 5.86
Color Thresholding None 86.45 37.32 35.26 95.14 21.65 21.42 41.77 1.92 1.87
MAESTER None 2.16 18.95 .94 22.08 40.30 16.64 0.29 5.03 0.27
SAM + IoU Tracking External 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
TriSAM External 84.12 66.75 59.28 78.41 74.97 62.14 31.35 25.57 16.39
Figure 5: Instance segmentation results of comparison methods on BvEM-Macaque. Different colors indicate different instances. Color thresholding and 3D UNet often produce false positives, whereas SAM + IoU tracking tends to miss a significant portion of blood vessels. Among the comparison methods, TriSAM segmentation stands out as the most effective.
4.2 Benchmark Results
Methods in Comparison. We compare the proposed TriSAM with both zero-shot baselines and supervised methods. The compared zero-shot methods are global color thresholding and SAM + IoU tracking. For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10x512x512). Then we label all voxels that are 3 standard derivation above mean as positive. Finally, connected components with less than 1000 voxels are filtered out. For SAM + IoU Tracking, we segment all objects in each z-slice of the dataset using automatic mask generation. Then we track each blood vessel using the first labelled slice as seeds. Our simple tracking algorithm finds the mask in next slice with maximum IoU with the current slice. If the max IoU is above a threshold, we assign this mask to the current object and continue tracking. We also tried SAM + IoU tracking with microSAM Archit et al. (2023b) weights that have been finetuned on EM images. This model however does not work well on our datasets. We expect this is because microSAM has been finetuned on high-resolution EM images and does not generalize to our low-resolution datasets. We further compare TriSAM with the supervised method 3D U-Net (The MICrONS Consortium et al., 2023). We use the implementation from Wolny et al. (2020). Individual models are trained on $5\%$ of each dataset. We use $64\times 192\times 192$ dimensional cubes for training and train for $20000$ iterations.
Results Analysis. The experimental results are shown in Table 2. First, both the Color Thresholding and SAM + IoU Tracking methods exhibit significant performance variability across various sub-datasets, highlighting the diversity of our dataset and the sensitivity of these methods to different species. Furthermore, both of these unsupervised methods demonstrate relatively poor performance, underscoring the challenges of the zero-shot setting in the BvEM dataset. Additionally, the 3D UNet, as a supervised learning approach, also yields very subpar results, indicating poor generalization of models trained with limited data. Finally, TriSAM outperforms the comparison methods by a large margin as it not only accurately segments the boundary but also tracks the blood vessels in the long term. Note that the results from the original dataset paper are also reported for reference.
Qualitative Results. The final instance segmentation results on BvEM-Macaque are shown in Figure 5. Color thresholding segments bright pixels, inadvertently capturing nuclei cells while overlooking darker pixels corresponding to blood vessels. Training the 3D UNet model with limited data results in confusion with background elements. IoU tracking fails to capture a significant portion of the blood vessel, revealing its ineffectiveness in tracking. TriSAM prediction emerges as the most accurate method, affirming the effectiveness of the proposed approach.
4.3 Ablation Study
We conducted a comprehensive series of ablation studies exclusively using the BvEM-Macaque dataset due to the computation constraints.
Table 3: Ablation study results on different plan selection strategies. The proposed Tri-plane approach achieves the best overall accuracy with comparable speed.
Method Precision (%) Recall (%) Accuracy (%) Speed (sec)
Single-Plane (z) 75.28 48.48 41.82 324
Single-Plane (y) 79.37 58.47 50.75 307
Single-Plane (x) 69.73 12.14 11.53 345
Single-Plane (fusion) 71.78 74.15 57.41 976
Tri-Plane 78.41 74.97 62.14 335
Figure 6: Comparison of segmentation results from one seed using different plane selection strategies.
Refer to captionRefer to caption
Figure 6: Comparison of segmentation results from one seed using different plane selection strategies.
Figure 7: Ablation study results on the hyperparameters $\eta$ and $\tau$.
Effect of the Tri-Plane Selection. We first compare our method with the single-plane methods to evaluate the effectiveness of Tri-Plane. For the single-plane method, we choose one plane as the main plane and only track along the chosen plane. The results on the BvEM-Macaque dataset are shown in Tabel 3. We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane. This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others. Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other. Instead, Tri-Plane exploits the blood vessel 3D structures by tracking the blood vessels along a suitable plane and attains the highest accuracy of 62.14%.
We visualize the segmentation results with one initial seed on BvEM-Macaque in Figure 7. We see that the performance is sensitive to the selection of the tracking plane. If the plane is not well-selected, the segmentation result can be empty as the example of tracking along the x-axis shows. The best result is tracking along the y-axis. However, it still falls short of the proposed method’s performance, as the latter takes into account potential turning points to exploit the 3D structure and perform long-term tracking.
Table 4: Ablation study results on different redirection strategies. The proposed recursive redirection approach achieves the best overall accuracy with comparable speed.
Strategy Precision (%) Recall (%) Accuracy (%) Speed (sec)
Naive 79.37 58.47 50.75 307
Dense Redirection 86.94 24.84 23.95 2001($\uparrow$ 552%)
Recursive Redirection 78.41 74.97 62.14 335 ($\uparrow$ 9%)
Effect of the Recursive Redirection Strategy. To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive redirection by considering potential turning points. To validate its effectiveness, we report the results on the BvEM-Macaque dataset in Table 4, where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included. One naive baseline is to remove the recursive redirection component and not consider any potential turning points. This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance. Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes. Unfortunately, it significantly increases the computation cost. However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse. This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
Table 5: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
Method Pre (%) Rec (%) Acc (%) Speed Parameters
Ours + SAM (Kirillov et al., 2023) 77.65 66.03 55.48 1535s 615M
Ours + MobileSAM (Zhang et al., 2023) 78.41 74.97 62.14 335s 9.66M
Effect of the Mobile-SAM Model. To further improve the efficiency, we use MobileSAM instead of the original SAM in practice. As shown in Table 5, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed. Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.
Quality Analysis We delve deeper into the plane dynamics in Figures 10 and 10. The majority of the selected plane predominantly tracks along the y-axis. This observation aligns with the experimental results presented in Table3, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis. This is because the blood vessel flows mainly along the y-axis as shown in Figure 7. In Figure 10, we explore the size dynamics using various methods. The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
Figure 8: Histogram of tracking planes.
Figure 9: Dynamics of tracking planes.
Figure 8: Histogram of tracking planes.
Figure 9: Dynamics of tracking planes.
Figure 10: Dynamics of size during tracking. (X) indicates the mean derivation is X.
5 Related Work
Blood Vessel Segmentation. Most existing VEM image segmentation algorithms were developed for neurons (Lee et al., 2019) and synapses (Wu et al., 2023a; Turner et al., 2020; Buhmann et al., 2021). The traditional VEM dataset size is too small to study blood vessel architecture. Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex (Shapson-Coe et al., 2021; The MICrONS Consortium et al., 2023). This marks a significant milestone in connectomics research. The blood vessels were segmented manually in the human cortex (Shapson-Coe et al., 2021) and automatically using fully convolutional neural network (FCN) Tetteh et al. (2020) or 3D U-Net (The MICrONS Consortium et al., 2023; Livne et al., 2019; Çiçek et al., 2016). Due to the burden of annotation, efforts have been made to decrease the need for annotations (Dang et al., 2022). For images acquired using other related imaging methods, such as light microscopy and MRI, a variety of methods were developed and covered by some reviews (Goni et al., 2022; Moccia et al., 2018). Briefly, the blood vessels could be enhanced based on the nature of tubeness using filters, such as diffusion filtering and ridge filtering. However, the segmentation of cortical blood vessels in VEM images is seldom explored, mainly because there is a lack of publicly available benchmarks.
Evaluation Metrics. Quantitative evaluation is normally conducted by comparing with manual annotations that are used as the ground truth. The evaluation metrics could be classified into two categories: pixel-based and skeleton-based (Moccia et al., 2018). Pixel-based metrics measure the difference between segmentation and ground truth at the pixel or voxel level, such as Accuracy and Dice Similarity Coefficient (Dice, 1945). Skeleton-based metrics measure the difference of skeletons or centerlines, such as Overlap and Overlap Until First Error (Metz et al., 2008). Skeleton-based metric requires the blood vessel skeleton that might not always be available. It also requires the computation of point-to-point correspondence between the ground truth and computed skeleltons. All existing evaluation metrics are developed using datasets acquired using other imaging methods and we did not find any specific metric for VEM datasets. Here we adopt the metrics Precision, Recall, and Accuracy as we are measuring instance-level pixel prediction.
Segment Anything-Based Models. As a foundation model for image segmentation, the recently proposed Segment Anything Model (Kirillov et al., 2023) has garnered significant attention. Due to its exceptional zero-shot performance, SAM has been extended to a variety of domains including object tracking (Cheng et al., 2023b; Yang et al., 2023), image inpainting (Yu et al., 2023), image mattting (Yao et al., 2023), super-resolution (Lu et al., 2023), 3D point cloud (Liu et al., 2023), and image editting (Gao et al., 2023). Despite SAM’s remarkable generalization capabilities, it still encounters several challenges in practical applications. One of these challenges is the huge computation costs due to the heavyweight image encoder. FastSAM (Zhao et al., 2023) adopted a conventional CNN detector with an instance segmentation branch for the segment anything task with real-time speed. MobileSAM (Zhang et al., 2023) proposed decoupled distillation to obtain a small image encoder, which achieved approximately five times faster speed compared to FastSAM while also being seven times smaller in size. Therefore the MobileSAM is employed in our proposed method. Another challenge lies in the unsatisfactory performance of SAM when confronted with special domains, such as medical (Chen et al., 2023) or biological (Archit et al., 2023a) images, particularly in the context of 3D data. Deng et al. (2023) assessed the SAM model’s zero-shot segmentation performance in the context of digital pathology and showed scenarios where SAM encounters difficulties. Mazurowski et al. (2023) extensively evaluates the SAM for medical image segmentation across 19 diverse datasets, highlighting SAM’s performance variability based on prompts and dataset characteristics. To address the domain gap between natural and medical images, SAM-Adapter (Chen et al., 2023), SAM-Med2D (Cheng et al., 2023a), and Medical SAM Adapter (Wu et al., 2023b) introduced Adapter modules and trained the Adapter with medical images. They attained good performance on various medical image segmentation tasks. MedSAM (Ma & Wang, 2023) adapted SAM with more than one million medical image-mask pairs and attained accurate segmentation results. MicroSAM (Archit et al., 2023a) also presented a segment anything model for microscopy by fine-tuning SAM with microscopy data. Differing from these approaches that require model fine-tuning for adaptation, our method lifts the blood vessel segmentation capabilities of SAM from 2D images to 3D volumes without any model fine-tuning.
6 Conclusion
In this paper, we have contributed the largest-to-date public benchmark, the BvEM dataset, for cortical blood vessel segmentation in 3D VEM images. By addressing a significant gap in the field of neuroimaging, we have laid the foundation for advancing the understanding of cerebral vasculature at the microscale and its intricate relationship with neural function. We also developed a zero-shot blood vessel segmentation method, TriSAM, based on the powerful SAM model, offering an efficient and accurate approach for segmenting blood vessels in VEM images. With Tri-Plane selection, SAM-based tracking, and recursive redirection, our TriSAM effectively exploits the 3D blood vessel structure and attains superior performance compared with existing zero-shot and supervised technologies on BvEM across three species, marking a critical step towards unlocking the mysteries of neurovascular coupling and its implications for brain health and pathology. With the availability of the BvEM dataset and the TriSAM method, researchers are now equipped with valuable tools to drive breakthroughs in VEM-based cortical blood vessel segmentation and further our understanding of the brain’s intricate vascular network.
Appendix A Appendix
A.1 Failure Cases
Despite our method’s significant progress, it still encounters challenges in some scenarios, as illustrated in Figure 11. The first issue is the seed problem, where the presence of a false positive seed can lead to erroneous predictions. Secondly, SAM could also fail to segment the mask. Lastly, tracking failure tends to occur at blood vessel conjunction points. These instances highlight the complexity of tracking in such intricate vascular networks and suggest the need for further refinements in our approach to address these specific failure cases effectively.
Figure 11: The failure cases of the proposed method. Three prediction and ground truth pairs from BvEM-Macaque are shown. Results are generated from one initial seed located at the origin point.
A.2 Algorithm
The algorithm of TriSAM is shown in Algorithm 1.
Algorithm 1 TriSAM algorithm
[…]
Initialize the turning point tree $TurningPointTree=s$
for ${\textnormal{t}}_{i}\quad in\quad\operatorname{UnsampledLeafNode}(TurningPointTree)$ do
[…]
     if $\operatorname{Valid}(TurningPoints)$ then
         $TurningPointTree=\operatorname{LeafSplit}(TurningPointTree,{\textnormal{t}}_{i},TurningPoints)$triangleright Grow the turning point tree
     end if
end for
[…]
Refer to captionRefer to captionRefer to caption
Figure 12: Dataset statistics: the histogram of the blood vessel radius.
A.3 Dataset statistics
The histogram of the blood vessels in the proposed dataset is shown in Figure 12.

### v2

@@
 Abstract
+While imaging techniques at macro and mesoscales have garnered substantial attention and resources, microscale VEM imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure.
 In this paper, we address a significant gap in the field of neuroimaging by introducing the largest-to-date public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in volume electron microscopy (VEM) images.
-The intricate relationship between cerebral blood vessels and neural function underscores the vital role of vascular analysis in understanding brain health.
-While imaging techniques at macro and mesoscales have garnered substantial attention and resources, the microscale VEM imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure.
-As researchers delve deeper into the microscale intricacies of cerebral vasculature, our BvEM benchmark represents a critical step toward unraveling the mysteries of neurovascular coupling and its impact on brain function and pathology.
-The BvEM dataset is based on VEM image volumes from three mammal species: adult mouse, macaque, and human.
+Our BvEM benchmark is based on VEM image volumes from three mammal species: adult mouse, macaque, and human.
 We standardized the resolution, addressed imaging variations, and meticulously annotated blood vessels through semi-automatic, manual, and quality control processes, ensuring high-quality 3D segmentation.
 Furthermore, we developed a zero-shot cortical blood vessel segmentation method named TriSAM, which leverages the powerful segmentation model SAM for 3D segmentation.
-To lift SAM from 2D segmentation to 3D volume segmentation, TriSAM employs a multi-seed tracking framework, leveraging the reliability of certain image planes for tracking while using others to identify potential turning points.
-This approach, consisting of Tri-Plane selection, SAM-based tracking, and recursive redirection, effectively achieves long-term 3D blood vessel segmentation without model training or fine-tuning.
+To extend SAM from 2D to 3D volume segmentation, TriSAM employs a multi-seed tracking framework, leveraging the reliability of certain image planes for tracking while using others to identify potential turning points.
+This approach effectively achieves long-term 3D blood vessel segmentation without model training or fine-tuning.
 Experimental results show that TriSAM achieved superior performances on the BvEM benchmark across three species.
-1 Introduction
+Index Terms:
+Blood vessel, Segment anything, Zero-shot segmentation
+Figure 1: Comparison of imaging modalities for blood vessel analysis.
+(a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details.
+(b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all the cell types and their cellular details.
+I Introduction
 With around 2% of body weight, our brain receives around 20% of blood supply.
-With inadequate blood supply, such as stroke, an average patient loses $\sim$1.9 million neurons (Saver, 2006).
-Alterations of blood vessel structures are observed in many brain diseases, such as Alzheimer’s and vascular dementia (Kalaria, 2010).
-Thus, blood vessels in the brain have been extensively investigated by a variety of imaging methods with different resolutions and structural details (Figure 1a).
-Compared to the macro-level imaging (e.g., CT and MRI (Dyer et al., 2017; McDonald & Choyke, 2003)) and mesoscale-level imaging (e.g.light microscopy Lochhead et al.
-(2023)), Volume Electron Microscopy (VEM) (Peddie & Collinson, 2014; Peddie et al., 2022) can reveal the detailed ultrastructure including endothelial cells, glial cells, and neurons (Figure 1b).
-Traditionally, the imaging methods at the macro and mesoscale are widely used and have produced a large amount of data, and a variety of image segmentation algorithms, public datasets, and evaluation methods have been developed  (Goni et al., 2022; Moccia et al., 2018).
-Recently, owing to the rapid improvement of imaging technology, the sample size of VEM is significantly increased covering all the layers of the cerebral cortex of mouse (The MICrONS Consortium et al., 2023)and human brain (Shapson-Coe et al., 2021), as well as the whole brain of fly (Dorkenwald et al., 2023).
-However, to robustly analyze blood vessels in VEM images, existing segmentation methods suffer from two major challenges: the diversity of the image appearance due to variations in the involved imaging pipeline including sample preparation, and the complexity of blood vessel morphology.
-Figure 1: Comparison of imaging modalities for blood vessel analysis.
-(a) Both microtomography ($\mu$CT) Dyer et al.
-(2017) and light microscopy (LM) Lochhead et al.
-(2023) can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details.
-(b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all the cell types and their details are visible for further investigations.
-To accelerate the method development, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images.
-The VEM image volumes of the BvEM dataset are from recent publications, which are the largest for each of the three mammal species: mouse, macaque, and human acquired at different VEM facilities.
+Most of the energy and nutrients are consumed by the neurons, and neuron function is sensitive to the blood supply [3, 4].
+The blood supply can even be adjusted following the consumption of brain regions, called neurovascular coupling [5].
+Alterations of blood vessel structures are observed in many brain diseases, e.g., Alzheimer’s and vascular dementia [6].
+Thus, blood vessels in the brain have been extensively investigated with various imaging modalities at different resolutions (Figure 1a).
+Compared to the macro-level imaging (e.g., CT [1] and MRI [7]) and mesoscale-level imaging (e.g., light microscopy [2]), volume electron microscopy (VEM) [8] can further reveal the detailed ultrastructure including all vascular cells (Figure 1b) for in-depth analysis.
+However, no large-scale annotated VEM dataset exists to develop and evaluate automated 3D blood vessel segmentation methods.
+Traditionally, the imaging methods at the macro and mesoscale are widely used and have produced a large amount of data, and a variety of image segmentation algorithms, public datasets, and evaluation methods have been developed  [9, 10].
+At the microscale level, the sample size of VEM is normally limited and most image analyses focus on neuron reconstruction, and blood vessels are largely ignored.
+Recently, owing to the rapid improvement of imaging technology, the sample size of VEM is significantly increased covering all the layers of the cerebral cortex of mouse [11]and human brain [12], as well as the whole brain of fly [13].
+Moreover, imaging the whole mouse brain using VEM technology is under planning [14].
+Thus, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images to foster segmentation method development.
+The raw image volumes are from recent publications, which are the largest for each of the three mammal species: mouse, macaque, and human acquired at different VEM facilities.
 We downsampled the volumes to a consistent resolution and performed extensive blood vessel annotation, including manual proofreading, semi-automatic segmentation error correction, and quality control, involving multiple rounds of scrutiny by neuroscience experts to ensure accuracy and completeness.
-We then propose a zero-shot 3D segmentation method named TriSAM based on the Segment Anything Model (SAM) (Kirillov et al., 2023) which is able to segment objects in the image given points or a bounding box as the input prompt.
-With a multi-seed tracking framework, our proposed TriSAM consists of three components: Tri-Plane selection, SAM-based tracking, and recursive redirection.
-Given the fact that tracking along blood vessel flow direction is easier, our Tri-Plane selection method chooses the best plane for tacking.
-To further exploit the 3D blood vessel structure and ensure long-term tracking, we propose to change the tracking direction at the potential turning points with a recursive redirection strategy.
-The proposed method achieves state-of-the-art performance compared to the baseline methods on the proposed BvEM benchmark across all three species.
-2 BvEM Dataset
-Dataset Design.
-To foster new methods tackling the challenges, we built our BvEM dataset on top of the largest publicly available VEM image volumes for samples from each of three mammal species: visual cortex from an adult mouse The MICrONS Consortium et al.
-(2023), superior temporal gyrus from an adult macaque Loomba et al.
-(2022), and temporal lobe from an adult human Shapson-Coe et al.
-(2021).
-Each dataset was acquired with different protocols at different VEM facilities and we refer readers to the respective papers for more details.
-Image Volume Standardization.
-We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and a manageable dataset size for computation and dissemination for AI researchers.
-Then, we trimmed off the rows and columns across all slices near the image boundary of the mouse and human data, where the blood vessels are hard to annotate due to the missing image content.
-We directly used the macaque dataset which was well-trimmed by its authors.
-As shown in Figure 2, the imaging quality and the appearance of blood vessels vary drastically across these three datasets, showcasing the cutting-edge large-scale VEM pipelines in the field.
-Figure 2: BvEM Dataset.
-We compiled the largest publicly available VEM dataset for each of the (a) mouse The MICrONS Consortium et al.
-(2023), macaque Loomba et al.
-(2022), and human Shapson-Coe et al.
-(2021) samples acquired by three different VEM labs.
-For each dataset, we downsampled the image volume to a similar resolution and trimmed off boundary regions.
-Then, we corrected the existing ground truth or instance segmentation results for each blood vessel instance (displayed in different colors).
+However, on the BvEM dataset, existing 3D blood vessel segmentation methods suffer from two major challenges: the diversity of the image appearance due to variations in the imaging pipeline and the complexity of the 3D blood vessel morphology.
+Conventional blood vessel segmentation heavily depends on a substantial volume of manually annotated data, a resource that is notably scarce in the existing literature.
+To address these challenges, we propose a zero-shot 3D segmentation method, TriSAM, leveraging the recent 2D segmentation foundation model, Segment Anything Model (SAM) [15], to handle the appearance diversity.
+To re-purpose SAM for 3D segmentation, we developed a multi-seed segment tracking framework based upon the video object segmentation method with SAM [16].
+As it is easier to track 2D blood vessel segments along the blood flow direction, TriSAM selects the best 2D plane for SAM-based tracking and further introduces a recursive seed sampling strategy that performs tri-plane selection at the potential turning points for efficiency.
+Experimental results demonstrate that the proposed TriSAM method significantly outperforms the prior state-of-the-art methods on the proposed BvEM benchmark across all three species.
+II Related Work
+Blood Vessel Segmentation.
+Most existing VEM image segmentation algorithms were developed for neurons [17] and synapses [18, 19, 20].
+The traditional VEM dataset size is too small to study blood vessel architecture.
+Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11].
+This marks a significant milestone in connectomics research.
+The blood vessels were segmented manually in the human cortex [12] and automatically using fully convolutional neural network (FCN) [21] or 3D U-Net [11, 22, 23].
+Due to the burden of annotation, efforts have been made to decrease the need for annotations [24].
+For images acquired using other related imaging methods, such as light microscopy and MRI, many methods were developed and covered by some reviews [9, 10].
+Briefly, the blood vessels could be enhanced based on the nature of tubeness using filters like diffusion and ridge filtering.
+Yet, the segmentation of cortical blood vessels in VEM images is seldom explored due to the absence of publicly available benchmarks.
+Evaluation Metrics..
+Quantitative evaluation is normally conducted by comparing with manual annotations that are used as the ground truth.
+The evaluation metrics could be classified into two categories: pixel-based and skeleton-based [10].
+Pixel-based metrics measure the difference between segmentation and ground truth at the pixel or voxel level, such as Accuracy and Dice Similarity Coefficient [25].
+Skeleton-based metrics measure the difference of skeletons or centerlines, such as Overlap and Overlap Until First Error [26].
+Skeleton-based metric requires the blood vessel skeleton that might not always be available.
+It also requires the computation of point-to-point correspondence between the ground truth and computed skeletons.
+All existing evaluation metrics are developed using datasets acquired using other imaging methods and we did not find any specific metric for VEM datasets.
+Most of them could be utilized in VEM images directly.
+Here we adopt the metrics Precision, Recall, and Accuracy as we measure instance-level pixel prediction.
+Segment Anything-Based Models.
+As a foundation model for image segmentation, the recently proposed Segment Anything Model [15] has garnered significant attention and has been extended to a variety of domains [27, 28, 29, 30] including object tracking [27, 16], image inpainting [28], image mattting [31], super-resolution [32], 3D point cloud [33], and image editting [34].
+Despite SAM’s remarkable generalization capabilities, it still encounters some challenges in practical applications, One of these challenges is the huge computation costs due to the heavyweight image encoder.
+FastSAM [35] adopted a conventional CNN detector with an instance segmentation branch for the segment anything task with real-time speed.
+MobileSAM [36] proposed decoupled distillation to obtain a small image encoder, which achieved approximately five times faster speed compared to FastSAM while also being seven times smaller in size.
+Therefore the MobileSAM is employed in our proposed method.
+One important challenge lies in the unsatisfactory performance of SAM when confronted with special domains, such as medical [37] or biological [38] images, particularly in the context of 3D data.
+Although SAM demonstrates impressive generalization capabilities, it faces challenges in specialized domains like medical [37, 39] or biological images [38], especially in handling 3D data.
+[30] assessed the SAM model’s zero-shot segmentation performance in the context of digital pathology and showed scenarios where SAM encounters difficulties.
+[40] extensively evaluates the SAM for medical image segmentation across 19 diverse datasets, highlighting SAM’s performance variability.
+To address the domain gap between natural and medical images, SAM-Adapter [37], SAM-Med2D [41], and Medical SAM Adapter [42] introduced Adapter modules and trained the Adapter with medical images.
+They attained good performance on various medical image segmentation tasks.
+MedSAM [29] adapted SAM with more than one million medical image-mask pairs and attained accurate segmentation results.
+MicroSAM [38] also presented a segment anything model for microscopy by fine-tuning SAM with microscopy data.
+Unlike these approaches that require model fine-tuning for adaptation, our method lifts the blood vessel segmentation capabilities of SAM from 2D images to 3D volumes without any model fine-tuning.
+Figure 2: The proposed BvEM dataset.
+We proofread the blood vessel instance segmentation (displayed in different colors) in the three largest publicly available VEM volumes: (a) mouse [11], macaque [43], and human [12] samples acquired at different VEM labs.
+III BvEM Dataset
+III-A Sample Information
+We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [^1] BvEM-Mouse can be visualized in: https://tinyurl.com/yc2s38x9 [11], superior temporal gyrus from an adult macaque [43], and temporal lobe from an adult human [12].
+Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details.
+TABLE I: Dataset information.
+Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined.
+The last column shows the max/total length of the blood vessel instances.
+Sample Resolution (nm) Size (voxel) Length: (mm)
+Mouse (TEM) 320$\times$ 256$\times$256 2495$\times$3571$\times$2495 1.6/1.7
+Macaque (SBEM) 240$\times$176$\times$176 450$\times$1271$\times$995 713.3/714.5
+Human (MultiSEM) 264$\times$256$\times$256 661$\times$7752$\times$13500 107.2/126.7
+III-B Image Volume Standardization
+We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and a manageable dataset size for AI researchers.
+Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content.
+As shown in Fig.
+2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
+III-C Blood Vessel Instance Segmentation Annotation
 […]
-Table 1: Dataset Statistics.
-Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined.
-Sample Microscope Resolution (nm) Volume size (voxel) Length: max/sum (mm)
-Mouse TEM 320$\times$ 256$\times$256 2495$\times$3571$\times$2495 1.6/1.7
-Macaque SBEM 240$\times$176$\times$176 450$\times$1271$\times$995 713.3/714.5
-Human MultiSEM 264$\times$256$\times$256 661$\times$7752$\times$13500 107.2/126.7
-Dataset Statistics.
-We summarize the dataset details in Table 1, where the dimension is in the “zyx" order.
-The BvEM-Macaque data has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human image data are around 80 and 121 times bigger than it.
-From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines with the Kimimaro software (Silversmith et al., 2021) and computed the length for each blood vessel instance.
-Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each dataset.
-Note that the BvEM-Human dataset is a thin slab, where many blood vessels are disconnected due to the limited spread.
-3 Method
-Conventional blood vessel segmentation heavily depends on a substantial volume of manually annotated data, a resource that is notably scarce in the existing literature.
-Therefore, we propose a zero-shot 3D blood vessel segmentation method based on SAM (Kirillov et al., 2023) which achieves superior performance without model training or fine-tuning.
+Figure 3: Dataset statistics: the histogram of the blood vessel radius.
+Figure 4: TriSAM framework.
+(a) Tri-plane selection is first proposed to select the best plane for tracking.
+(b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking direction.
+(c) Recursive seed sampling exploits potential turning points for long-term tracking.
+III-D Dataset Statistics
+We summarize the dataset details in Tab.
+I.
+The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively.
+From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [45] and computed the length for each blood vessel instance.
+Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume.
+The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
+IV Method
+IV-A Overview
+We formulate the 3D blood vessel segmentation task as a multi-seed video object segmentation task: generating one 3D segment per seed through tracking 2D SAM results along one fixed direction and the final output is the connected component of the combined segmentation results (Fig.
+5a).
+However, unlike videos, the 3D blood vessel segment can be tracked from different axes, whereas 2D SAM segmentation results can be of drastically different qualities.
+As shown in Fig.
+5b, the seed generation can have many errors leading to incomplete segmentation results.
+To achieve long-term tracking, the framework needs a recursive seed sampling module to propose potential turning points as new seeds based on segmentation results.
+As shown in Fig.
+5c, the ground truth segment on some 2D planes can have such an unnatural shape that SAM fails to segment it, not to mention its drastic shape change along the axis.
+Ideally, one axis is along the blood flow direction where the 2D tracking plane is an ellipse shape, e.g., zx plane in the example.
+Thus, it is essential to have a tri-plane selection module in the framework.
+The proposed TriSAM framework has the following three modules: tri-plane selection, SAM-based tracking, and recursive seed sampling (Fig. 4).
 Figure 5: Problem formulation and challenges.
@@
 There are two main challenges for such an approach: (b) generating high-quality initial seeds (blue dots) and minimizing the false positive (red crosses) and false negative errors (yellow circles); (c) adaptively selecting tracking directions as the quality of SAM segmentation results can vary significantly.
-3.1 Problem formulation: 3D Segmentation as multi-seed tracking
-To apply SAM to 3D VEM images, a naive method is to combine SAM with a tracking algorithm.
-However, blood vessels are sensitive to split error requiring robust long-term tracking which is a remaining challenge in object tracking literature.
-Another possible solution is to use a 3D adaptor with 2D SAM features which requires a large amount of labeled data for training.
-However, labeled training data is limited since the annotation for VEM images needs the involvement of experts.
-To this end, we formulate 3D segmentation as a multi-seed tracking framework shown in Figure 3a that doesn’t require any annotated data or robust long-term tracking algorithms.
-To achieve long-term tracking with multi-seed tracking, we leverage the fact that one of the planes is reliable for tracking while other planes indicate potential turning points.
-In particular, different from videos, the VEM data can be interpolated from different planes (e.g.
-yx/zy/zx planes).
-Tracking along some planes is extremely difficult since the shape and scale change dramatically.
-Therefore, it is essential to select which plane should be used for tracking.
-To this end, we propose a Tri-Plane Selection method to select the best plane for tracking based on the potential blood flow direction.
-To achieve long-term tracking of blood vessels, a Recursive Redirection method is proposed to keep tracking at potential turning points.
-Based on these two components, a simple SAM-based Tracking is effective in achieving long-term 3D blood vessel segmentation.
-3.2 TriSAM 3D Segmentation
-The proposed TriSAM 3D segmentation method has three components: Tri-Plane selection, SAM-based tracking, and recursive redirection.
-The initial seeds are first generated automatically or manually.
-Then, TriSAM 3D segmentation is performed on each seed and the segmentation results are combined as the final segmentation.
-The pipeline of the proposed method is shown in Figure 4.
-Initial Seed Generation.
+IV-B Tri-Plane Selection
+Rather than videos, 3D VEM images can be interpreted in different planes.
+In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically.
+This phenomenon can be further confirmed by the experiment shown in Table III in which tracking along the $y$ axis is more effective than tracking along the $x$ axis.
+Given the seed position as a prompt, we run SAM on each of the tri-planes of the 3D image volume centered at the seed.
+Note that the plane that is most tangent to the blood flow direction has the smallest segment size, capturing the blood vessel’s cross-section (Fig.
+4a, xy-plane).
+In addition, SAM outputs the probability for the segment result, reflecting the confidence of the naturalness of its shape.
+Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$.
+TriSAM terminates when no plane is selected.
+IV-C SAM-based Tracking
+Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis.
+If we use the center of the current segment as the point prompt for the next slice, SAM may output segments with inconsistent sizes along the axis due to the lack of size information in the point prompt.
+If we use the bounding box of the current segment as the box prompt, SAM may not track in the long term as the output enclosed in the bounding box is getting smaller and smaller.
+Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results.
+TriSAM terminates when SAM returns no prediction for the next image slice.
+IV-D Recursive Seed Sampling
+The above modules work well when the blood vessel travels in one direction.
+In reality, blood vessels not only turn in different directions but also bifurcate, leading to drastic shape and scale changes as shown in Fig.
+11.
+Thus, we need a module to propose potential turning points to recursively run TriSAM, linking short-term segment tracks into long-term segment tracks.
+SAM-based tracking is simple and fast for short-term tracking.
+But it may fail when shape and scale change dramatically as shown in Fig.
+11.
+Therefore, it relies on the large number of seeds which is hard to generate as shown in Fig.
+5(b).
+To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels.
+Given the 3D segment from the short-term track after TriSAM terminates, we first use the Farthest Point Sampling (FPS) to find potential endpoints.
+Then, we add these points to the global seed tree whose root note is the input seed.
+To avoid repeating the track along the same direction as the previous short-term track, these newly sampled seeds store the previous tracking direction.
+For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions.
+During each iteration, we take an unvisited node in the seed tree, mark it as visited, and run TriSAM along a different tracking direction.
+With such a seed tree data structure, we run TriSAM recursively until all potential points are visited.
+For each endpoint, we re-run the above two TriSAM modules to track segments along directions that are different from the current tracking axis.
+Mathematically, we denote the 3D coordinate of a turning point as ($x_{t}$,$y_{t}$,$z_{t}$).
+At this point, we perform the SAM operation on three planes separately.
+We use $M_{z}(x,y,z_{t})$, $M_{y}(x,y_{t},z)$, and $M_{x}(x_{t},y,z)$ to represent the segmented mask region along xy plane, xz plane, and yz plane, respectively.
+Assuming that the chosen best plane is xz plane which means we will track along the y-axis, we will obtain new turning points based on the segmented masks on the non-best planes $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$.
+Since the boundary region of the mask corresponds to where blood vessels disappear, it indicates locations where there is a change in the direction of blood vessel flow.
+So, the boundaries in the y-direction of both $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$ indicate locations where blood vessels change direction along the y-axis.
+We further define the maximum and minimum values of the y-coordinate within the $M_{z}(x,y,z_{t})$ region as $y_{max}^{z_{t}}$ and $y_{min}^{z_{t}}$, respectively.
+Similarly, we can define $y_{max}^{x_{t}}$ and $y_{min}^{x_{t}}$.
+Then the new turning points can be defined as $\{(x,y,z)|(x,y,z)\in M_{z}(x,y_{max}^{z_{t}},z_{t})\cup M_{z}(x,y_{min}^{z_{t}},z_{t})\cup M_{x}(x_{t},y_{max}^{x_{t}},z)\cup M_{x}(x_{t},y_{min}^{x_{t}},z)\}$.
+The turning point will be dropped if it hits the volume boundary or the probability of its segment is less than the threshold $\tau$.
+After updating the turning point tree, we then sample a new leaf node for the next iteration until all leaf nodes are sampled.
+In practice, we maintain a tree of seed points where the root is the initial seed.
+In each iteration, we sample a left node in the tree, mark it as visited, and run TriSAM.
+Empirically, we find the non-best planes potentially identify turning points or endpoints accurately.
+Therefore, we propose to generate the turning points according to the segmentation results at non-best planes.
+Potential turning points are identified as points in these segments with the smallest and largest z-values.
+The algorithm of TriSAM is shown in Algorithm 1.
+Algorithm 1 TriSAM algorithm
+[…]
+[…]
+Initialize the seed tree $SeedTree=s$
+for ${\textnormal{t}}_{i}\quad in\quad\operatorname{RemainNode}(SeedTree)$ do
+[…]
+[…]
+[…]
+[…]
+     if $\operatorname{Valid}(Seeds)$ then
+         $SeedTree=\operatorname{LeafSplit}(SeedTree,{\textnormal{t}}_{i},Seeds)$triangleright Grow the seed tree
+     end if
+end for
+[…]
+V Experiments
+In this section, we present TriSAM’s experimental results. We begin with settings and implementation details. Then we show the benchmark results. Finally, we perform ablations for further analysis.
+V-A Experimental Settings
+Implementation Details.
+Since the volume is too large, non-overlapping subvolumes are used for segmentation and the results are fused to form the final prediction.
 Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background.
@@
 To improve efficiency, we only keep the center of each connected component as the final seeds.
-In practice, the threshold is set to $\eta I_{m}$ where $I_{m}$ is the maximum intensity of the volume.
-Tri-Plane Selection.
-Rather than videos, 3D VEM images can be interpreted in different planes.
-In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically.
-As shown in Figure 4, tracking along the bottom plane is hard because the shape is complex and changes quickly.
-This phenomenon can be further confirmed by the experiment shown in Figure 7 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis.
-Tracking along the plane according to the blood flow direction is ideal and this can be achieved based on choosing a plane with smaller segmentation as shown in Figure 4.
-Furthermore, we take into account the probability associated with the SAM segmentation result, as it reflects the confidence level in identifying blood vessels.
-Overall, the plane with the smallest segments and a probability higher than the threshold $\tau$ is selected as the best plane.
-SAM-based Tracking.
-By tracking along the optimally selected plane, changes in shape and size of the blood vessel are more stable compared to other planes.
-Therefore, we use the enlarged bounding box and the center of the segment as the prompts for the next slice.
-In particular, we first pad the bounding box of the segment as the bounding box prompt and compute the center of the segment as the point prompt.
-These prompts are used for segmenting blood vessels in the next slice.
-The tracking will stop when the probability of the segment is less than a threshold $\tau$.
-Figure 4: Left: SAM-based tracking.
-Right: our proposed TriSAM.
-(a) Tri-Plane selection is first proposed to select the best plane for tracking.
-(b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking direction.
-(c) Recursive redirection exploits potential turning points for long-term tracking.
-Recursive Redirection.
-SAM-based tracking is simple and fast for short-term tracking.
-But it may fail when shape and scale change dramatically as shown in Figure 3c.
-To enable long-term tracking, we propose a recursive redirection method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels.
-Specifically, for each seed, we maintain a turning point tree where the root is the initial seed.
-In each iteration, we sample a left node in the turning point tree and perform Tri-Plane selection and track along the chosen plane.
-As blood vessels may change directions within this iteration, we then split this node to obtain new potential turning points.
-We find that the non-best planes potentially identify turning points or endpoints accurately as shown in Figure 4, Therefore, we propose to generate the turning points according to the segmentation results at non-best planes.
-In particular, when tracking along the z-axis, we apply SAM to segment blood vessels in the zx and zy planes.
-Potential turning points are identified as points in these segments with the smallest and largest z-values.
-4 Experiments
-In this section, we present TriSAM’s experimental results. We begin with settings and implementation details. Then we show the benchmark results. Finally, we perform ablations for further analysis.
-4.1 Experimental Settings and Implementation Details
-Dataset Split.
-We used the whole dataset for evaluation.
-For the supervised setting, we randomly sampled subvolumes are annotated.
-We sampled 4 subvolumes of size 256$\times$256$\times$256 for the BvEM-Macaque dataset amounting to around 10% of the dataset, and 12 subvolumes of size 384$\times$384$\times$384 for each of the BvEM-Mouse and BvEM-Human datasets amounting to around 1% of each dataset.
-These annotated subvolumes were divided into a 3-1 train-valid split.
-Implementation Details.
-Since the volume is too large, non-overlapping subvolumes are used for segmentation and the results are fused to form the final prediction.
 For BvEM-Macaque, we directly predict the whole volume.
-For others, $k\times 1024\times 1024$ subvolumes are used since $1024\times 1024$ is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse.
-For preprocessing, we utilize defliker method and temporal smoothing on the z-axis to remove the artifacts and missing slices.
+For others, $k\times 1024\times 1024$ subvolumes are used since it is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse.
+For preprocessing, we utilize deflickering method and temporal smoothing on the z-axis to remove the artifacts and missing slices.
 For postprocessing, holes are filled and small connected components are removed after the segmentation.
-Global color thresholding is used to generate the initial seeds and $\eta$ is set to 0.98 as shown Figure 7 (left).
-Given the significantly greater complexity of human blood vessels, we employ the manually generated seeds for the challenging BvEM-Human dataset.
-For SAM-based tracking, the probability threshold $\tau$ is set to 0.8 according to the experiment shown in Figure 7 (right).
-Unless explicitly stated otherwise, we use MobileSAM (Zhang et al., 2023) instead of the standard SAM (Kirillov et al., 2023) in all conducted experiments to improve the inference speed.
+Global color thresholding is used to generate the initial seeds and $\eta$ is set to 0.98 as shown Fig.
+8 (left).
+Given the significantly greater complexity of human blood vessels, we employ the manually generated seeds for the challenging BvEM-Human volume.
+For SAM-based tracking, the probability threshold $\tau$ is set to 0.8.
+according to the experiment shown in Fig.
+8 (right).
+Unless explicitly stated otherwise, we use MobileSAM [36] instead of the standard SAM [15] in all conducted experiments to improve the inference speed.
 All experiments are conducted on an NVIDIA-A100 GPU.
-Evaluation Metrics. Following Weigert et al. (2020), we use Precision, Recall, and Accuracy as the metrics to evaluate the performance:
-$Precision=\frac{TP}{TP+FP},\quad Recall=\frac{TP}{TP+FN},\quad Accuracy=\frac{TP}{TP+FP+FN},$ (1)
+Evaluation Metrics. We use the Precision, Recall, and Accuracy metric defined in [46], where $Accuracy=\frac{TP}{TP+FP+FN}$.
+$Precision$ $=\frac{TP}{TP+FP},$ (1)
+$Recall$ $=\frac{TP}{TP+FN},$ (2)
+$Accuracy$ $=\frac{TP}{TP+FP+FN},$ (3)
 where $TP$, $FP$, and $FN$ are instance-level true positive, false positive, and false negative respectively.
@@
 In particular, the Hungarian algorithm is used to match ground-truth instances and predicted instances with negative Accuracy as the cost matrix.
-Table 2: Benchmark results on the proposed BvEM dataset.
-We evaluate the initial blood vessel annotation to show the amount of proofreading effort.
-We compare the proposed TriSAM with various baseline methods under fully supervised with limited training data, unsupervised, and zero-shot settings.
-Method Training Data BvEM-Mouse BvEM-Macaque BvEM-Human
+We used the whole dataset for evaluation and computed the score on the largest instance segment of each volume due to its dominating size.
+TABLE II: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the significant amount of our proofreading effort.
+Method Setting BvEM-Mouse BvEM-Macaque BvEM-Human
 Pre Rec Acc Pre Rec Acc Pre Rec Acc
-Initial Annotation  [^1] Results from original dataset papers.
-BvEM-Mouse and BvEM-Macaque are automatically segmented while BvEM-Human is manually labeled.
-Unknown 93.74 36.62 35.74 1.65 23.17 1.57 100.00 25.68 25.68
-3D UNet Limited 13.45 0.91 0.67 16.04 89.65 15.75 68.63 2.46 2.43
-nnUNET Limited 3.54 49.55 6.59 24.34 29.57 23.74 3.44 23.20 5.86
-Color Thresholding None 86.45 37.32 35.26 95.14 21.65 21.42 41.77 1.92 1.87
-MAESTER None 2.16 18.95 .94 22.08 40.30 16.64 0.29 5.03 0.27
-SAM + IoU Tracking External 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
-TriSAM External 84.12 66.75 59.28 78.41 74.97 62.14 31.35 25.57 16.39
-Figure 5: Instance segmentation results of comparison methods on BvEM-Macaque.
+Initial annotation N/A 93.74 36.62 35.74 1.65 23.17 1.57 100.00 25.68 25.68
+Color thresholding N/A 86.45 37.32 35.26 95.14 21.65 21.42 41.77 1.92 1.87
+MAESTER [47] unsupervised 2.16 18.95 0.94 22.08 40.30 16.64 0.29 5.03 0.27
+3D UNet [11] supervised 13.45 0.91 0.67 16.56 86.95 16.16 68.63 2.46 2.43
+nnUNET [48] supervised 3.54 49.55 6.59 24.34 29.57 23.74 3.44 23.20 5.86
+SAM+IoU tracking zero-shot 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
+TriSAM (ours) zero-shot 84.12 66.75 59.28 78.41 74.97 62.14 31.35 25.57 16.39
+Figure 6: Qualitative instance segmentation results on the BvEM-Macaque volume.
 Different colors indicate different instances.
@@
 Among the comparison methods, TriSAM segmentation stands out as the most effective.
-4.2 Benchmark Results
+V-B Benchmark Results
 Methods in Comparison.
@@
 For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10x512x512).
-Then we label all voxels that are 3 standard derivation above mean as positive.
+Then we label all voxels that are 3 standard derivations above mean as positive.
 Finally, connected components with less than 1000 voxels are filtered out.
 For SAM+IoU Tracking, we segment all objects in each z-slice of the dataset using automatic mask generation.
-Then we track each blood vessel using the first labelled slice as seeds.
-Our simple tracking algorithm finds the mask in next slice with maximum IoU with the current slice.
+Then we track each blood vessel using the first labeled slice as seeds.
+Our simple tracking algorithm finds the mask in the next slice with maximum IoU with the current slice.
 If the max IoU is above a threshold, we assign this mask to the current object and continue tracking.
-We also tried SAM + IoU tracking with microSAM Archit et al.
-(2023b) weights that have been finetuned on EM images.
-This model however does not work well on our datasets.
-We expect this is because microSAM has been finetuned on high-resolution EM images and does not generalize to our low-resolution datasets.
-We further compare TriSAM with the supervised method 3D U-Net (The MICrONS Consortium et al., 2023).
-We use the implementation from Wolny et al.
-(2020).
-Individual models are trained on $5\%$ of each dataset.
-We use $64\times 192\times 192$ dimensional cubes for training and train for $20000$ iterations.
+We also tried SAM+IoU tracking with microSAM [49] weights that have been finetuned on EM images.
+This model however does not work well on our dataset.
+We expect this is because microSAM has been finetuned on high-resolution EM images and does not generalize to our low-resolution dataset.
+We further compare TriSAM with the supervised method 3D U-Net [11].
+We use the implementation from [50].
+For supervised methods, we cropped subvolumes at the center of each volume composing approximately 10% in size.
+These annotated subvolumes were divided into a 1-1 train-val split.
 Results Analysis.
-The experimental results are shown in Table 2.
-First, both the Color Thresholding and SAM + IoU Tracking methods exhibit significant performance variability across various sub-datasets, highlighting the diversity of our dataset and the sensitivity of these methods to different species.
+The results are shown in Tab.
+II.
+First, both the Color Thresholding and SAM+IoU Tracking methods exhibit significant performance variability across three volumes, highlighting the diversity of our dataset and the sensitivity of these methods to different species.
 Furthermore, both of these unsupervised methods demonstrate relatively poor performance, underscoring the challenges of the zero-shot setting in the BvEM dataset.
-Additionally, the 3D UNet, as a supervised learning approach, also yields very subpar results, indicating poor generalization of models trained with limited data.
-Finally, TriSAM outperforms the comparison methods by a large margin as it not only accurately segments the boundary but also tracks the blood vessels in the long term.
-Note that the results from the original dataset paper are also reported for reference.
+Additionally, the 3D UNet, as a supervised learning approach, also yields subpar results, indicating poor generalization of models trained with limited data.
+Finally, TriSAM significantly outperforms other methods as it not only accurately segments the boundary but also tracks the blood vessels in the long term.
 Qualitative Results.
-The final instance segmentation results on BvEM-Macaque are shown in Figure 5.
+The final instance segmentation results on BvEM-Macaque are shown in Fig.
+6.
 Color thresholding segments bright pixels, inadvertently capturing nuclei cells while overlooking darker pixels corresponding to blood vessels.
@@
 IoU tracking fails to capture a significant portion of the blood vessel, revealing its ineffectiveness in tracking.
-TriSAM prediction emerges as the most accurate method, affirming the effectiveness of the proposed approach.
-4.3 Ablation Study
+TriSAM prediction emerges as the most accurate method, affirming its effectiveness.
+V-C Ablation Studies
 We conducted a comprehensive series of ablation studies exclusively using the BvEM-Macaque dataset due to the computation constraints.
-Table 3: Ablation study results on different plan selection strategies. The proposed Tri-plane approach achieves the best overall accuracy with comparable speed.
-Method Precision (%) Recall (%) Accuracy (%) Speed (sec)
+Effect of the Tri-plane Selection.
+We first compare our method with the single-plane methods to evaluate the effectiveness of the tri-plane.
+For the single-plane method, we choose one plane as the main plane and only track along the chosen plane.
+The results on the BvEM-Macaque volume are shown in Tab.
+III.
+We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane.
+This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others.
+Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other.
+Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.
+TABLE III: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
+Method Pre. (%) Rec. (%) Acc. (%) Speed (sec)
 Single-plane (z) 75.28 48.48 41.82 324
@@
 Tri-plane 78.41 74.97 62.14 335
-Figure 6: Comparison of segmentation results from one seed using different plane selection strategies.
+Figure 7: Segmentation results with different plane selection strategies.
 Refer to captionRefer to caption
-Figure 6: Comparison of segmentation results from one seed using different plane selection strategies.
-Figure 7: Ablation study results on the hyperparameters $\eta$ and $\tau$.
-Effect of the Tri-Plane Selection.
-We first compare our method with the single-plane methods to evaluate the effectiveness of Tri-Plane.
-For the single-plane method, we choose one plane as the main plane and only track along the chosen plane.
-The results on the BvEM-Macaque dataset are shown in Tabel 3.
-We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane.
-This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others.
-Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other.
-Instead, Tri-Plane exploits the blood vessel 3D structures by tracking the blood vessels along a suitable plane and attains the highest accuracy of 62.14%.
-We visualize the segmentation results with one initial seed on BvEM-Macaque in Figure 7.
+Figure 7: Segmentation results with different plane selection strategies.
+Figure 8: Ablation study results on hyperparameters $\eta$ and $\tau$.
+Figure 9: Histogram of the tri-plane selection.
+Figure 10: Dynamics of the tri-plane selection.
+Figure 9: Histogram of the tri-plane selection.
+Figure 10: Dynamics of the tri-plane selection.
+Figure 11: Dynamics of segment size.
+We visualize the segmentation results with one initial seed on BvEM-Macaque in Fig.
+8.
+The origin point is the seed location.
 We see that the performance is sensitive to the selection of the tracking plane.
 If the plane is not well-selected, the segmentation result can be empty as the example of tracking along the x-axis shows.
+The best result is tracking along the y-axis which is still worse than the proposed method since it considers potential turning points to leverage the 3D structure.
 The best result is tracking along the y-axis.
 However, it still falls short of the proposed method’s performance, as the latter takes into account potential turning points to exploit the 3D structure and perform long-term tracking.
-Table 4: Ablation study results on different redirection strategies. The proposed recursive redirection approach achieves the best overall accuracy with comparable speed.
-Strategy Precision (%) Recall (%) Accuracy (%) Speed (sec)
+TABLE IV: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
+Strategy Pre. (%) Rec. (%) Acc. (%) Speed (sec)
 Naive 79.37 58.47 50.75 307
-Dense Redirection 86.94 24.84 23.95 2001($\uparrow$ 552%)
-Recursive Redirection 78.41 74.97 62.14 335 ($\uparrow$ 9%)
+Dense sampling 86.94 24.84 23.95 2001($\uparrow$ 552%)
+Recursive sampling 78.41 74.97 62.14 335 ($\uparrow$ 9%)
 Effect of the Recursive Redirection Strategy.
-To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive redirection by considering potential turning points.
-To validate its effectiveness, we report the results on the BvEM-Macaque dataset in Table 4, where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included.
-One naive baseline is to remove the recursive redirection component and not consider any potential turning points.
+To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive seed sampling by considering potential turning points.
+To validate its effectiveness, we report the results on the BvEM-Macaque volume in Tab.
+IV where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included.
+Compared to the baseline without recursive seed sampling “naive” and dense seed sampling “dense”, the proposed method achieves the best performance with less than 10% running time increase.
+One naive baseline is to remove the recursive seed sampling component and not consider any potential turning points.
 This strategy is simple and fast but it fails to exploit the 3D shape prior leading to poor performance.
 Another strategy is to select the best plane for every tracking step/slice, which densely performs SAM segmentation on each step across three planes.
+Therefore, the performance is limited.
+The proposed method achieves the best performance with less than 10% running time increase on the BvEM-Macaque dataset.
 Unfortunately, it significantly increases the computation cost.
+Compared to our method, DenseSAM’s running time is 5.97 times longer since it needs to segment 3 planes for every tracking step.
 However, we were surprised to observe that the performance of the Dense Redirection strategy was even worse.
 This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
-Table 5: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
-Method Pre (%) Rec (%) Acc (%) Speed Parameters
-Ours + SAM (Kirillov et al., 2023) 77.65 66.03 55.48 1535s 615M
-Ours + MobileSAM (Zhang et al., 2023) 78.41 74.97 62.14 335s 9.66M
-Effect of the Mobile-SAM Model.
+TABLE V: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
+Backbone Pre (%) Rec (%) Acc (%) Speed Parameters
+SAM [15] 77.65 66.03 55.48 1535s 615M
+MobileSAM [36] 78.41 74.97 62.14 335s 9.66M
+Quality Analysis.
+We delve deeper into the plane dynamics in Fig.
+11 and 11.
+The majority of the selected plane predominantly tracks along the y-axis.
+This observation aligns with the experimental results presented in Tab.III, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis.
+This is because the blood vessel flows mainly along the y-axis in the test volume.
+In Fig.
+11, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses.
+The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
+Other Ablations.
 To further improve the efficiency, we use MobileSAM instead of the original SAM in practice.
-As shown in Table 5, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed.
-Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.
-Quality Analysis We delve deeper into the plane dynamics in Figures 10 and 10.
-The majority of the selected plane predominantly tracks along the y-axis.
-This observation aligns with the experimental results presented in Table3, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis.
-This is because the blood vessel flows mainly along the y-axis as shown in Figure 7.
-In Figure 10, we explore the size dynamics using various methods.
-The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
-Figure 8: Histogram of tracking planes.
-Figure 9: Dynamics of tracking planes.
-Figure 8: Histogram of tracking planes.
-Figure 9: Dynamics of tracking planes.
-Figure 10: Dynamics of size during tracking. (X) indicates the mean derivation is X.
-5 Related Work
-Blood Vessel Segmentation.
-Most existing VEM image segmentation algorithms were developed for neurons (Lee et al., 2019) and synapses (Wu et al., 2023a; Turner et al., 2020; Buhmann et al., 2021).
-The traditional VEM dataset size is too small to study blood vessel architecture.
-Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex (Shapson-Coe et al., 2021; The MICrONS Consortium et al., 2023).
-This marks a significant milestone in connectomics research.
-The blood vessels were segmented manually in the human cortex (Shapson-Coe et al., 2021) and automatically using fully convolutional neural network (FCN) Tetteh et al.
-(2020) or 3D U-Net (The MICrONS Consortium et al., 2023; Livne et al., 2019; Çiçek et al., 2016).
-Due to the burden of annotation, efforts have been made to decrease the need for annotations (Dang et al., 2022).
-For images acquired using other related imaging methods, such as light microscopy and MRI, a variety of methods were developed and covered by some reviews (Goni et al., 2022; Moccia et al., 2018).
-Briefly, the blood vessels could be enhanced based on the nature of tubeness using filters, such as diffusion filtering and ridge filtering.
-However, the segmentation of cortical blood vessels in VEM images is seldom explored, mainly because there is a lack of publicly available benchmarks.
-Evaluation Metrics.
-Quantitative evaluation is normally conducted by comparing with manual annotations that are used as the ground truth.
-The evaluation metrics could be classified into two categories: pixel-based and skeleton-based (Moccia et al., 2018).
-Pixel-based metrics measure the difference between segmentation and ground truth at the pixel or voxel level, such as Accuracy and Dice Similarity Coefficient (Dice, 1945).
-Skeleton-based metrics measure the difference of skeletons or centerlines, such as Overlap and Overlap Until First Error (Metz et al., 2008).
-Skeleton-based metric requires the blood vessel skeleton that might not always be available.
-It also requires the computation of point-to-point correspondence between the ground truth and computed skeleltons.
-All existing evaluation metrics are developed using datasets acquired using other imaging methods and we did not find any specific metric for VEM datasets.
-Here we adopt the metrics Precision, Recall, and Accuracy as we are measuring instance-level pixel prediction.
-Segment Anything-Based Models.
-As a foundation model for image segmentation, the recently proposed Segment Anything Model (Kirillov et al., 2023) has garnered significant attention.
-Due to its exceptional zero-shot performance, SAM has been extended to a variety of domains including object tracking (Cheng et al., 2023b; Yang et al., 2023), image inpainting (Yu et al., 2023), image mattting (Yao et al., 2023), super-resolution (Lu et al., 2023), 3D point cloud (Liu et al., 2023), and image editting (Gao et al., 2023).
-Despite SAM’s remarkable generalization capabilities, it still encounters several challenges in practical applications.
-One of these challenges is the huge computation costs due to the heavyweight image encoder.
-FastSAM (Zhao et al., 2023) adopted a conventional CNN detector with an instance segmentation branch for the segment anything task with real-time speed.
-MobileSAM (Zhang et al., 2023) proposed decoupled distillation to obtain a small image encoder, which achieved approximately five times faster speed compared to FastSAM while also being seven times smaller in size.
-Therefore the MobileSAM is employed in our proposed method.
-Another challenge lies in the unsatisfactory performance of SAM when confronted with special domains, such as medical (Chen et al., 2023) or biological (Archit et al., 2023a) images, particularly in the context of 3D data.
-Deng et al.
-(2023) assessed the SAM model’s zero-shot segmentation performance in the context of digital pathology and showed scenarios where SAM encounters difficulties.
-Mazurowski et al.
-(2023) extensively evaluates the SAM for medical image segmentation across 19 diverse datasets, highlighting SAM’s performance variability based on prompts and dataset characteristics.
-To address the domain gap between natural and medical images, SAM-Adapter (Chen et al., 2023), SAM-Med2D (Cheng et al., 2023a), and Medical SAM Adapter (Wu et al., 2023b) introduced Adapter modules and trained the Adapter with medical images.
-They attained good performance on various medical image segmentation tasks.
-MedSAM (Ma & Wang, 2023) adapted SAM with more than one million medical image-mask pairs and attained accurate segmentation results.
-MicroSAM (Archit et al., 2023a) also presented a segment anything model for microscopy by fine-tuning SAM with microscopy data.
-Differing from these approaches that require model fine-tuning for adaptation, our method lifts the blood vessel segmentation capabilities of SAM from 2D images to 3D volumes without any model fine-tuning.
-6 Conclusion
-In this paper, we have contributed the largest-to-date public benchmark, the BvEM dataset, for cortical blood vessel segmentation in 3D VEM images.
-By addressing a significant gap in the field of neuroimaging, we have laid the foundation for advancing the understanding of cerebral vasculature at the microscale and its intricate relationship with neural function.
-We also developed a zero-shot blood vessel segmentation method, TriSAM, based on the powerful SAM model, offering an efficient and accurate approach for segmenting blood vessels in VEM images.
-With Tri-Plane selection, SAM-based tracking, and recursive redirection, our TriSAM effectively exploits the 3D blood vessel structure and attains superior performance compared with existing zero-shot and supervised technologies on BvEM across three species, marking a critical step towards unlocking the mysteries of neurovascular coupling and its implications for brain health and pathology.
-With the availability of the BvEM dataset and the TriSAM method, researchers are now equipped with valuable tools to drive breakthroughs in VEM-based cortical blood vessel segmentation and further our understanding of the brain’s intricate vascular network.
-Appendix A Appendix
-A.1 Failure Cases
+As shown in Table V, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed.
+Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.Fig.
+8 illustrates the performance with varying hyperparameters $\eta$ and $\tau$.
+Figure 12: The failure cases of the proposed method. Three prediction and ground truth pairs are shown.
+V-D Failure Cases
 Despite our method’s significant progress, it still encounters challenges in some scenarios, as illustrated in Figure 12.
@@
 These instances highlight the complexity of tracking in such intricate vascular networks and suggest the need for further refinements in our approach to address these specific failure cases effectively.
-Figure 11: The failure cases of the proposed method. Three prediction and ground truth pairs from BvEM-Macaque are shown. Results are generated from one initial seed located at the origin point.
-A.2 Algorithm
-The algorithm of TriSAM is shown in Algorithm 1.
-Algorithm 1 TriSAM algorithm
-[…]
-[…]
-Initialize the turning point tree $TurningPointTree=s$
-for ${\textnormal{t}}_{i}\quad in\quad\operatorname{UnsampledLeafNode}(TurningPointTree)$ do
-[…]
-[…]
-[…]
-[…]
-     if $\operatorname{Valid}(TurningPoints)$ then
-         $TurningPointTree=\operatorname{LeafSplit}(TurningPointTree,{\textnormal{t}}_{i},TurningPoints)$triangleright Grow the turning point tree
-     end if
-end for
-[…]
-Refer to captionRefer to captionRefer to caption
-Figure 12: Dataset statistics: the histogram of the blood vessel radius.
-A.3 Dataset statistics
-The histogram of the blood vessels in the proposed dataset is shown in Figure 12.
+VI Conclusion
+In this paper, we contribute the largest-to-date public benchmark, the BvEM dataset, for cortical blood vessel segmentation in 3D VEM images.
+We also developed a zero-shot blood vessel segmentation method, TriSAM, based on the powerful SAM model, offering an efficient and accurate approach for segmenting blood vessels in VEM images.
+With tri-plane selection, SAM-based tracking, and recursive seed sampling, our TriSAM effectively exploits the 3D blood vessel structure and attains superior performance compared with existing zero-shot and supervised technologies on BvEM across three species, marking a critical step towards unlocking the mysteries of neurovascular coupling and its implications for brain health and pathology.
+With the availability of the BvEM dataset and the TriSAM method, researchers are now equipped with valuable tools to drive breakthroughs in VEM-based cortical blood vessel segmentation and further our understanding of the brain’s intricate vascular network.
+By addressing a significant gap in the field of neuroimaging, we have laid the foundation for advancing the understanding of cerebral vasculature at the microscale and its intricate relationship with neural function.

### v3

@@
 Abstract
-While imaging techniques at macro and mesoscales have garnered substantial attention and resources, microscale VEM imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure.
-In this paper, we address a significant gap in the field of neuroimaging by introducing the largest-to-date public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in volume electron microscopy (VEM) images.
-Our BvEM benchmark is based on VEM image volumes from three mammal species: adult mouse, macaque, and human.
+While imaging techniques at macro and mesoscales have garnered substantial attention and resources, microscale Volume Electron Microscopy (vEM) imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure.
+In this paper, we address a significant gap in this field of neuroimaging by introducing the first-in-class public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in vEM images.
+Our BvEM benchmark is based on vEM image volumes from three mammals: adult mouse, macaque, and human.
 We standardized the resolution, addressed imaging variations, and meticulously annotated blood vessels through semi-automatic, manual, and quality control processes, ensuring high-quality 3D segmentation.
@@
 Experimental results show that TriSAM achieved superior performances on the BvEM benchmark across three species.
-Index Terms:
+Our dataset, code, and model are available online at https://jia-wan.github.io/bvem.
+{IEEEkeywords}
 Blood vessel, Segment anything, Zero-shot segmentation
@@
 (a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details.
-(b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all the cell types and their cellular details.
-I Introduction
+(b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all types of cells with details.
+1 Introduction
 With around 2% of body weight, our brain receives around 20% of blood supply.
@@
 Thus, blood vessels in the brain have been extensively investigated with various imaging modalities at different resolutions (Figure 1a).
-Compared to the macro-level imaging (e.g., CT [1] and MRI [7]) and mesoscale-level imaging (e.g., light microscopy [2]), volume electron microscopy (VEM) [8] can further reveal the detailed ultrastructure including all vascular cells (Figure 1b) for in-depth analysis.
+Compared to the macro-level imaging (e.g., CT [1] , MRI [7]) and mesoscale-level imaging (e.g., light microscopy [2]), volume electron microscopy (VEM) [8] can further reveal the detailed ultrastructure including all vascular cells (Figure 1b) for in-depth analysis.
 However, no large-scale annotated VEM dataset exists to develop and evaluate automated 3D blood vessel segmentation methods.
@@
 Thus, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images to foster segmentation method development.
-The raw image volumes are from recent publications, which are the largest for each of the three mammal species: mouse, macaque, and human acquired at different VEM facilities.
+The raw image volumes are from recent publications, which are the largest for each of the three mammals: mouse, macaque, and human acquired at different VEM facilities.
 We downsampled the volumes to a consistent resolution and performed extensive blood vessel annotation, including manual proofreading, semi-automatic segmentation error correction, and quality control, involving multiple rounds of scrutiny by neuroscience experts to ensure accuracy and completeness.
@@
 As it is easier to track 2D blood vessel segments along the blood flow direction, TriSAM selects the best 2D plane for SAM-based tracking and further introduces a recursive seed sampling strategy that performs tri-plane selection at the potential turning points for efficiency.
+The proposed method is similar to Flood-Filling Networks (FFN) [17] by extending the mask in 3D space.
+Instead of training a 3D neural network, we utilize SAM for prediction since it generalize better.
+Second, FFN densely extends to 3D while the proposed method extends along the blood vessels.
+Finally, FFN uses only one seed while we are using growing seeds for long-term tracking.
 Experimental results demonstrate that the proposed TriSAM method significantly outperforms the prior state-of-the-art methods on the proposed BvEM benchmark across all three species.
-II Related Work
-Blood Vessel Segmentation.
+2 Related Work
+Blood vessel segmentation methods.
 Most existing VEM image segmentation algorithms were developed for neurons [18] and synapses [19, 20, 21].
@@
 For images acquired using other related imaging methods, such as light microscopy and MRI, many methods were developed and covered by some reviews [9, 10].
-Briefly, the blood vessels could be enhanced based on the nature of tubeness using filters like diffusion and ridge filtering.
+Briefly, the blood vessels could be enhanced based on the tubular nature using filters like diffusion and ridge filtering.
 Yet, the segmentation of cortical blood vessels in VEM images is seldom explored due to the absence of publicly available benchmarks.
-Evaluation Metrics..
-Quantitative evaluation is normally conducted by comparing with manual annotations that are used as the ground truth.
-The evaluation metrics could be classified into two categories: pixel-based and skeleton-based [10].
-Pixel-based metrics measure the difference between segmentation and ground truth at the pixel or voxel level, such as Accuracy and Dice Similarity Coefficient [25].
-Skeleton-based metrics measure the difference of skeletons or centerlines, such as Overlap and Overlap Until First Error [26].
-Skeleton-based metric requires the blood vessel skeleton that might not always be available.
-It also requires the computation of point-to-point correspondence between the ground truth and computed skeletons.
-All existing evaluation metrics are developed using datasets acquired using other imaging methods and we did not find any specific metric for VEM datasets.
-Most of them could be utilized in VEM images directly.
-Here we adopt the metrics Precision, Recall, and Accuracy as we measure instance-level pixel prediction.
-Segment Anything-Based Models.
+Segment anything-based methods.
 As a foundation model for image segmentation, the recently proposed Segment Anything Model [15] has garnered significant attention and has been extended to a variety of domains [26, 27, 28, 29] including object tracking [26, 16], image inpainting [27], image mattting [30], super-resolution [31], 3D point cloud [32], and image editting [33].
@@
 We proofread the blood vessel instance segmentation (displayed in different colors) in the three largest publicly available VEM volumes: (a) mouse [11], macaque [42], and human [12] samples acquired at different VEM labs.
-III BvEM Dataset
-III-A Sample Information
-We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [^1] BvEM-Mouse can be visualized in: https://tinyurl.com/yc2s38x9 [11], superior temporal gyrus from an adult macaque [43], and temporal lobe from an adult human [12].
+3 BvEM Dataset
+3.1 Dataset Description
+Sample information.
+We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [11], superior temporal gyrus from an adult macaque [42], and temporal lobe from an adult human [12].
 Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details.
-TABLE I: Dataset information.
+Table 1: Dataset information.
 Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined.
@@
 Human (MultiSEM) 264$\times$256$\times$256 661$\times$7752$\times$13500 107.2/126.7
-III-B Image Volume Standardization
-We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and a manageable dataset size for AI researchers.
+Resolution standardization.
+We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and the nature of capillary diameter.
 Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content.
@@
 2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
-III-C Blood Vessel Instance Segmentation Annotation
+3.2 Dataset Annotation
 […]
+Semi-automatic segmentation error correction.
+We used the VAST lite software [43] which can accelerate the proofreading process by using provided segmentation results as the drawing or filling template instead of delineating segments manually.
+Quality control.
+We used 3D visualizations and skeleton analysis to detect remaining segmentation errors.
+Each volume was proofread by two neuroscience experts in multiple rounds until no disagreement.
 Figure 3: Dataset statistics: the histogram of the blood vessel radius.
+3.3 Dataset Statistics
+We summarize the dataset details in Tab.
+1.
+The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively.
+From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [44] and computed the length for each blood vessel instance.
+Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume.
+The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
+Figure 4: Problem formulation and challenges.
+(a) The 3D segmentation task can be formulated as the multi-seed object tracking problem.
+There are two main challenges for such an approach: (b) generating high-quality initial seeds (blue dots) and minimizing the false positive (red crosses) and false negative errors (yellow circles); (c) adaptively selecting tracking directions as the quality of SAM segmentation results can vary significantly.
 Figure 5: TriSAM framework.
@@
 (c) Recursive seed sampling exploits potential turning points for long-term tracking.
-III-D Dataset Statistics
-We summarize the dataset details in Tab.
-I.
-The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively.
-From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [45] and computed the length for each blood vessel instance.
-Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume.
-The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
-IV Method
-IV-A Overview
+4 Method
+4.1 Overview
 We formulate the 3D blood vessel segmentation task as a multi-seed video object segmentation task: generating one 3D segment per seed through tracking 2D SAM results along one fixed direction and the final output is the connected component of the combined segmentation results (Fig.
-5a).
+4a).
 However, unlike videos, the 3D blood vessel segment can be tracked from different axes, whereas 2D SAM segmentation results can be of drastically different qualities.
 As shown in Fig.
-5b, the seed generation can have many errors leading to incomplete segmentation results.
+4b, the seed generation can have many errors leading to incomplete segmentation results.
 To achieve long-term tracking, the framework needs a recursive seed sampling module to propose potential turning points as new seeds based on segmentation results.
 As shown in Fig.
-5c, the ground truth segment on some 2D planes can have such an unnatural shape that SAM fails to segment it, not to mention its drastic shape change along the axis.
+4c, the ground truth segment on some 2D planes can have such an unnatural shape that SAM fails to segment it, not to mention its drastic shape change along the axis.
 Ideally, one axis is along the blood flow direction where the 2D tracking plane is an ellipse shape, e.g., zx plane in the example.
@@
 The proposed TriSAM framework has the following three modules: tri-plane selection, SAM-based tracking, and recursive seed sampling (Fig. 5).
-Figure 5: Problem formulation and challenges.
-(a) The 3D segmentation task can be formulated as the multi-seed object tracking problem.
-There are two main challenges for such an approach: (b) generating high-quality initial seeds (blue dots) and minimizing the false positive (red crosses) and false negative errors (yellow circles); (c) adaptively selecting tracking directions as the quality of SAM segmentation results can vary significantly.
-IV-B Tri-Plane Selection
+4.2 Tri-Plane Selection
 Rather than videos, 3D VEM images can be interpreted in different planes.
 In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically.
-This phenomenon can be further confirmed by the experiment shown in Table III in which tracking along the $y$ axis is more effective than tracking along the $x$ axis.
+This phenomenon can be further confirmed by the experiment shown in Table 3 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis.
 Given the seed position as a prompt, we run SAM on each of the tri-planes of the 3D image volume centered at the seed.
 Note that the plane that is most tangent to the blood flow direction has the smallest segment size, capturing the blood vessel’s cross-section (Fig.
-4a, xy-plane).
+5a, xy-plane).
 In addition, SAM outputs the probability for the segment result, reflecting the confidence of the naturalness of its shape.
-Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$.
+Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$ (Select in Alg.
+1).
 TriSAM terminates when no plane is selected.
-IV-C SAM-based Tracking
-Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis.
+4.3 SAM-based Tracking
+Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis (SAMBasedTrack in Alg.
+1).
 If we use the center of the current segment as the point prompt for the next slice, SAM may output segments with inconsistent sizes along the axis due to the lack of size information in the point prompt.
@@
 Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results.
-TriSAM terminates when SAM returns no prediction for the next image slice.
-IV-D Recursive Seed Sampling
+TriSAM terminates when SAM returns no prediction for the next image slice (probability $<\tau$ ).
+4.4 Recursive Seed Sampling
 The above modules work well when the blood vessel travels in one direction.
@@
 Therefore, it relies on the large number of seeds which is hard to generate as shown in Fig.
-5(b).
-To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels.
+4(b).
+To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels (RecursiveSample in Alg.
+1).
 Given the 3D segment from the short-term track after TriSAM terminates, we first use the Farthest Point Sampling (FPS) to find potential endpoints.
-Then, we add these points to the global seed tree whose root note is the input seed.
+Then, we add these points to the global seed tree whose root note is the input seed (LeafSplit in Alg.
+1).
 To avoid repeating the track along the same direction as the previous short-term track, these newly sampled seeds store the previous tracking direction.
-For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions.
+For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions (Valid in Alg.
+1).
 During each iteration, we take an unvisited node in the seed tree, mark it as visited, and run TriSAM along a different tracking direction.
@@
 […]
-Initialize the seed tree $SeedTree=s$
+Initialize the seed tree $SeedTree=\{s\}$
 for ${\textnormal{t}}_{i}\quad in\quad\operatorname{RemainNode}(SeedTree)$ do
@@
 […]
-V Experiments
-In this section, we present TriSAM’s experimental results. We begin with settings and implementation details. Then we show the benchmark results. Finally, we perform ablations for further analysis.
-V-A Experimental Settings
+5 Experiments
+5.1 Experimental Settings
 Implementation Details.
@@
 We used the whole dataset for evaluation and computed the score on the largest instance segment of each volume due to its dominating size.
-TABLE II: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the significant amount of our proofreading effort.
+Table 2: Benchmark results on the proposed BvEM dataset. We evaluate the initial blood vessel annotation to show the significant amount of our proofreading effort.
 Method Setting BvEM-Mouse BvEM-Macaque BvEM-Human
@@
 nnUNET [47] supervised 3.54 49.55 6.59 24.34 29.57 23.74 3.44 23.20 5.86
-SAM+IoU tracking zero-shot 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
+SAM+IoU tracking[15] zero-shot 63.59 0.27 0.27 74.39 1.89 1.88 18.19 23.58 12.92
 TriSAM (ours) zero-shot 84.12 66.75 59.28 78.41 74.97 62.14 31.35 25.57 16.39
@@
 Among the comparison methods, TriSAM segmentation stands out as the most effective.
-V-B Benchmark Results
+5.2 Benchmark Results
 Methods in Comparison.
@@
 The results are shown in Tab.
-II.
+2.
 First, both the Color Thresholding and SAM+IoU Tracking methods exhibit significant performance variability across three volumes, highlighting the diversity of our dataset and the sensitivity of these methods to different species.
@@
 TriSAM prediction emerges as the most accurate method, affirming its effectiveness.
-V-C Ablation Studies
+5.3 Ablation Studies
 We conducted a comprehensive series of ablation studies exclusively using the BvEM-Macaque dataset due to the computation constraints.
@@
 The results on the BvEM-Macaque volume are shown in Tab.
-III.
+3.
 We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane.
@@
 Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.
-TABLE III: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
+Table 3: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
 Method Pre. (%) Rec. (%) Acc. (%) Speed (sec)
@@
 Figure 7: Segmentation results with different plane selection strategies.
-Refer to captionRefer to caption
-Figure 7: Segmentation results with different plane selection strategies.
 Figure 8: Ablation study results on hyperparameters $\eta$ and $\tau$.
-Figure 9: Histogram of the tri-plane selection.
-Figure 10: Dynamics of the tri-plane selection.
 Figure 9: Histogram of the tri-plane selection.
@@
 However, it still falls short of the proposed method’s performance, as the latter takes into account potential turning points to exploit the 3D structure and perform long-term tracking.
-TABLE IV: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
+Table 4: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
 Strategy Pre. (%) Rec. (%) Acc. (%) Speed (sec)
@@
 To validate its effectiveness, we report the results on the BvEM-Macaque volume in Tab.
-IV where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included.
+4 where the runtime comparison for segmentation prediction on the entire BvEM-Macaque data is also included.
 Compared to the baseline without recursive seed sampling “naive” and dense seed sampling “dense”, the proposed method achieves the best performance with less than 10% running time increase.
@@
 This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
-TABLE V: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
+Table 5: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
 Backbone Pre (%) Rec (%) Acc (%) Speed Parameters
@@
 The majority of the selected plane predominantly tracks along the y-axis.
-This observation aligns with the experimental results presented in Tab.III, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis.
+This observation aligns with the experimental results presented in Tab.3, where it is evident that a single plane tracking with the y-axis outperforms the z-axis and x-axis.
 This is because the blood vessel flows mainly along the y-axis in the test volume.
@@
 To further improve the efficiency, we use MobileSAM instead of the original SAM in practice.
-As shown in Table V, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed.
+As shown in Table 5, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed.
 Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.Fig.
@@
 Figure 12: The failure cases of the proposed method. Three prediction and ground truth pairs are shown.
-V-D Failure Cases
+5.4 Failure Cases
 Despite our method’s significant progress, it still encounters challenges in some scenarios, as illustrated in Figure 12.
@@
 These instances highlight the complexity of tracking in such intricate vascular networks and suggest the need for further refinements in our approach to address these specific failure cases effectively.
-VI Conclusion
+6 Conclusion
 In this paper, we contribute the largest-to-date public benchmark, the BvEM dataset, for cortical blood vessel segmentation in 3D VEM images.

### v4

@@
 Blood vessel, Segment anything, Zero-shot segmentation
-Figure 1: Comparison of imaging modalities for blood vessel analysis.
-(a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details.
-(b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all types of cells with details.
+Figure 1: Imaging modalities for blood vessel analysis.
+(a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can only capture blood vessels in the cortex at the sub-micron resolution without ultrastructure details.
+(b) Volume electron microscopy (VEM) can show unbiased details of the vasculature including all types of cells at a higher resolution.
 1 Introduction
@@
 The traditional VEM dataset size is too small to study blood vessel architecture.
-Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11].
-This marks a significant milestone in connectomics research.
-The blood vessels were segmented manually in the human cortex [12] and automatically using fully convolutional neural network (FCN) [22] or 3D U-Net [11, 23, 24].
+Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11], providing rich details of vasculature structure at scale.
+However, due to the lack of efficient and effective automatic methods, the blood vessels were segmented manually in the human cortex dataset [12].
+On another front, automatic blood vessel segmentation methods have been developed for other image modalities, such as light microscopy and MRI, at a lower resolution.
+We refer to  [9, 10] for detailed reviews on filter-based methods [22, 23] and deep learning methods [24, 25].
 Due to the burden of annotation, efforts have been made to decrease the need for annotations [26].
-For images acquired using other related imaging methods, such as light microscopy and MRI, many methods were developed and covered by some reviews [9, 10].
-Briefly, the blood vessels could be enhanced based on the tubular nature using filters like diffusion and ridge filtering.
-Yet, the segmentation of cortical blood vessels in VEM images is seldom explored due to the absence of publicly available benchmarks.
+In this paper, we propose the first VEM cortical blood vessel benchmark to foster novel methods.
 Segment anything-based methods.
@@
 3.1 Dataset Description
-Sample information.
+Tissue samples.
 We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [11], superior temporal gyrus from an adult macaque [43], and temporal lobe from an adult human [12].
 Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details.
+As shown in Fig.
+2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
+Image volumes.
+We processed the original image volumes into a standardized form that is suitable for benchmarking.
+We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and the nature of capillary diameter.
+Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content.
+3.2 Dataset Annotation
+[…]
+Automatic error detection.
+For the false split errors, which are the majority source of error, we computed the skeleton of the segmentation and detected all the skeleton endpoints that do not touch the volume boundary as candidates.
+For the false merge errors, we computed the intersection-over-union (IoU) of 2D segments on neighboring slices and detected areas of small IoU value as candidates.
+Manual proofreading.
+We used the VAST lite software [44] to accelerate the manual proofreading process by using the provided dense segmentation results as templates.
+Instead of manually delineating segments manually, proofreaders can coarsely draw or fill the segment mask that is snapped to the detailed boundary in the template.
+Assisted with the 3D blood vessel visualizations, two neuroscience experts proofread each volume in multiple rounds until no disagreement.
 Table 1: Dataset information.
@@
 Human (MultiSEM) 264$\times$256$\times$256 661$\times$7752$\times$13500 107.2/126.7
-Resolution standardization.
-We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and the nature of capillary diameter.
-Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content.
-As shown in Fig.
-2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
-3.2 Dataset Annotation
-[…]
-Semi-automatic segmentation error correction.
-We used the VAST lite software [43] which can accelerate the proofreading process by using provided segmentation results as the drawing or filling template instead of delineating segments manually.
-Quality control.
-We used 3D visualizations and skeleton analysis to detect remaining segmentation errors.
-Each volume was proofread by two neuroscience experts in multiple rounds until no disagreement.
-Figure 3: Dataset statistics: the histogram of the blood vessel radius.
 3.3 Dataset Statistics
-We summarize the dataset details in Tab.
-1.
-The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively.
+As shown in Tab.
+1, the BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively.
 From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [45] and computed the length for each blood vessel instance.
 Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume.
-The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
-Figure 4: Problem formulation and challenges.
-(a) The 3D segmentation task can be formulated as the multi-seed object tracking problem.
-There are two main challenges for such an approach: (b) generating high-quality initial seeds (blue dots) and minimizing the false positive (red crosses) and false negative errors (yellow circles); (c) adaptively selecting tracking directions as the quality of SAM segmentation results can vary significantly.
+The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3.
+The mouse and the human dataset have similar radius distributions peaked around capillaries, while the macaque volume has no capillaries.
+Figure 3: Dataset statistics. Empirical distribution of blood vessel radius.
 Figure 4: TriSAM framework.
 (a) Tri-plane selection is first proposed to select the best plane for tracking.
-(b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking direction.
+(b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking axis.
 (c) Recursive seed sampling exploits potential turning points for long-term tracking.
@@
 4.1 Overview
-We formulate the 3D blood vessel segmentation task as a multi-seed video object segmentation task: generating one 3D segment per seed through tracking 2D SAM results along one fixed direction and the final output is the connected component of the combined segmentation results (Fig.
-4a).
-However, unlike videos, the 3D blood vessel segment can be tracked from different axes, whereas 2D SAM segmentation results can be of drastically different qualities.
-As shown in Fig.
-4b, the seed generation can have many errors leading to incomplete segmentation results.
-To achieve long-term tracking, the framework needs a recursive seed sampling module to propose potential turning points as new seeds based on segmentation results.
-As shown in Fig.
-4c, the ground truth segment on some 2D planes can have such an unnatural shape that SAM fails to segment it, not to mention its drastic shape change along the axis.
-Ideally, one axis is along the blood flow direction where the 2D tracking plane is an ellipse shape, e.g., zx plane in the example.
-Thus, it is essential to have a tri-plane selection module in the framework.
-The proposed TriSAM framework has the following three modules: tri-plane selection, SAM-based tracking, and recursive seed sampling (Fig. 5).
-4.2 Tri-Plane Selection
-Rather than videos, 3D VEM images can be interpreted in different planes.
-In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically.
-This phenomenon can be further confirmed by the experiment shown in Table 3 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis.
-Given the seed position as a prompt, we run SAM on each of the tri-planes of the 3D image volume centered at the seed.
-Note that the plane that is most tangent to the blood flow direction has the smallest segment size, capturing the blood vessel’s cross-section (Fig.
-5a, xy-plane).
-In addition, SAM outputs the probability for the segment result, reflecting the confidence of the naturalness of its shape.
-Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$ (Select in Alg.
+We aim to leverage the 2D generalist Segment Anything Model (SAM) to build a zero-shot 3D VEM blood vessel segmentation method, which can be widely applied to various image volumes.
+Given an input seed position, existing SAM-based tracking methods work well when the blood vessel travels only in one axis.
+However, in reality, blood vessels not only turn in alternating directions but also bifurcate, leading to drastic shape and scale changes.
+Instead, we represent the whole connected blood vessel segment that contains the input seed as a graph, $G=(V,E)$, where each node $v\in V$ is a tubular branch segment that does not bifurcate.
+Thus, the segmentation problem can be solved with the graph traversal algorithm.
+After visiting one node (i.e., running SAM-based tracking methods given the seed), we sample new seed positions to find neighboring nodes to segment recursively.
+Thus, the proposed TriSAM framework has the following three modules: (1) Tri-plane selection module to pick the best 2D plane to segment the object (Fig.
+4a); (2) SAM-based tracking module to predict 2D masks along the selected axis (Fig.
+4b); (3) Turning point sampling module to propose potential turning points as new seeds based on existing segmentation results (Fig.
+4c) to recursively grow all branches of the blood vessel.
+Note that the second module can be any existing SAM-based tracking method.
+4.2 TriSAM Framework
+We first explain the three modules in the TriSAM framework (Fig.
+4) and then provide the algorithmic description (Alg.
 1).
-TriSAM terminates when no plane is selected.
-4.3 SAM-based Tracking
-Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis (SAMBasedTrack in Alg.
+In practice, we sample many initial seeds, run the TriSAM framework for each seed, and fuse the segmentation results.
+Tri-plane selection module.
+Unlike videos, 3D VEM image volumes $I$ can be segmented on 2D planes along different axes, $p\in\{x,y,z\}$, based on the object morphology.
+However, tracking along some axis $p$ can be harder than others due to the irregular shape of the 2D cross-section of the 3D blood vessels on the corresponding 2D plane $I_{p}$.
+We later verify such a phenomenon empirically as shown in Table 3 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis for a given seed position.
+This module aims to select the best axis $p^{*}$ for tracking, given the SAM segmentation results, $\mathrm{SAM}(I_{p},s)$, on each tri-plane $I_{p}$ centered at the seed position $s$ as the approximation of the blood vessel cross-sections.
+In theory, the ideal axis for tracking is along the blood flow direction, where the corresponding 2D image plane is tangent to the blood flow direction with the smallest cross-section segment area (Fig.
+4a, xy-plane).
+In addition, SAM outputs the probability, $P_{\mathrm{SAM}}$, for the segment result, reflecting the confidence in the naturalness of its shape.
+Thus, to combine the two cues above, we pick the axis with the smallest segment size on the 2D plane and a probability of at least the threshold $\tau$ (plane-select in Alg.
 1).
-If we use the center of the current segment as the point prompt for the next slice, SAM may output segments with inconsistent sizes along the axis due to the lack of size information in the point prompt.
-If we use the bounding box of the current segment as the box prompt, SAM may not track in the long term as the output enclosed in the bounding box is getting smaller and smaller.
-Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results.
-TriSAM terminates when SAM returns no prediction for the next image slice (probability $<\tau$ ).
-4.4 Recursive Seed Sampling
-The above modules work well when the blood vessel travels in one direction.
-In reality, blood vessels not only turn in different directions but also bifurcate, leading to drastic shape and scale changes as shown in Fig.
-11.
-Thus, we need a module to propose potential turning points to recursively run TriSAM, linking short-term segment tracks into long-term segment tracks.
-SAM-based tracking is simple and fast for short-term tracking.
-But it may fail when shape and scale change dramatically as shown in Fig.
-11.
-Therefore, it relies on the large number of seeds which is hard to generate as shown in Fig.
-4(b).
-To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels (RecursiveSample in Alg.
+$p^{*}=\arg\min_{p}\mathrm{Area}(\mathrm{SAM}(I_{p},s)),~{}\mathrm{s.t.}~{}P_{\mathrm{SAM}}(I_{p},s)>\tau.$ (1)
+To avoid an infinite recursion, traditional graph traversal algorithms check if a node is marked visited.
+In our case, a tubular branch segment may grow into different blood vessel segments along different tracking axes, e.g., bifurcation regions.
+Thus, we mark the combination of the seed and the tri-plane tracking axis $(s,p^{*})$ visited if seed $s$ falls in the segment that is predicted through tracking along $p^{*}$ axis (Visited in Alg.
 1).
-Given the 3D segment from the short-term track after TriSAM terminates, we first use the Farthest Point Sampling (FPS) to find potential endpoints.
-Then, we add these points to the global seed tree whose root note is the input seed (LeafSplit in Alg.
+Note that $p^{*}=\phi$ if no plane is selected and Visited($(s,p^{*})$) is True.
+To implement the “Visited” function, we store both the segmentation result and the $p^{*}$, which is omitted in Alg.
+1 for simplicity.
+SAM-based tracking module.
+Given the seed position $s$, the selected tracking axis $p^{*}$ and SAM’s initial 2D segment mask, we need to produce a 3D segment by tracking the 2D mask with SAM in both directions (e.g., $x+$ and $x-$).
+Note that naively propagating the mask center or bounding box as the SAM prompt for the next image slice leads to poor segmentation results, as SAM may output segments with inconsistent sizes or shrinking sizes along the axis respectively.
+Many existing sophisticated SAM-based tracking methods can be directly used.
+This module aims to provide a simple yet effective approach by generating better SAM prompts for the next image slice.
+Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results (SAM-Track in Alg.
 1).
-To avoid repeating the track along the same direction as the previous short-term track, these newly sampled seeds store the previous tracking direction.
-For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions (Valid in Alg.
-1).
-During each iteration, we take an unvisited node in the seed tree, mark it as visited, and run TriSAM along a different tracking direction.
-With such a seed tree data structure, we run TriSAM recursively until all potential points are visited.
-For each endpoint, we re-run the above two TriSAM modules to track segments along directions that are different from the current tracking axis.
-Mathematically, we denote the 3D coordinate of a turning point as ($x_{t}$,$y_{t}$,$z_{t}$).
-At this point, we perform the SAM operation on three planes separately.
-We use $M_{z}(x,y,z_{t})$, $M_{y}(x,y_{t},z)$, and $M_{x}(x_{t},y,z)$ to represent the segmented mask region along xy plane, xz plane, and yz plane, respectively.
-Assuming that the chosen best plane is xz plane which means we will track along the y-axis, we will obtain new turning points based on the segmented masks on the non-best planes $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$.
-Since the boundary region of the mask corresponds to where blood vessels disappear, it indicates locations where there is a change in the direction of blood vessel flow.
-So, the boundaries in the y-direction of both $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$ indicate locations where blood vessels change direction along the y-axis.
-We further define the maximum and minimum values of the y-coordinate within the $M_{z}(x,y,z_{t})$ region as $y_{max}^{z_{t}}$ and $y_{min}^{z_{t}}$, respectively.
-Similarly, we can define $y_{max}^{x_{t}}$ and $y_{min}^{x_{t}}$.
-Then the new turning points can be defined as $\{(x,y,z)|(x,y,z)\in M_{z}(x,y_{max}^{z_{t}},z_{t})\cup M_{z}(x,y_{min}^{z_{t}},z_{t})\cup M_{x}(x_{t},y_{max}^{x_{t}},z)\cup M_{x}(x_{t},y_{min}^{x_{t}},z)\}$.
-The turning point will be dropped if it hits the volume boundary or the probability of its segment is less than the threshold $\tau$.
-After updating the turning point tree, we then sample a new leaf node for the next iteration until all leaf nodes are sampled.
-In practice, we maintain a tree of seed points where the root is the initial seed.
-In each iteration, we sample a left node in the tree, mark it as visited, and run TriSAM.
-Empirically, we find the non-best planes potentially identify turning points or endpoints accurately.
-Therefore, we propose to generate the turning points according to the segmentation results at non-best planes.
-Potential turning points are identified as points in these segments with the smallest and largest z-values.
-The algorithm of TriSAM is shown in Algorithm 1.
-Algorithm 1 TriSAM algorithm
-[…]
-[…]
-Initialize the seed tree $SeedTree=\{s\}$
-for ${\textnormal{t}}_{i}\quad in\quad\operatorname{RemainNode}(SeedTree)$ do
+$prompt=\{(x,y),(x,y,\gamma w,\gamma h)\},$ (2)
+where $(x,y,\gamma w,\gamma h)$ is the bounding box of the segmentation and $\gamma$ is the scaling factor.
+To ensure the quality of the predicted SAM mask, the proposed tracking module terminates when the SAM probability for the predicted mask is lower than the threshold $\tau$.
+Turning point sampling module.
+As discussed above, SAM-based tracking module can not handle blood vessel segments that change directions or bifurcate, where the 2D segment along the original direction changes drastically.
+Thus, when the SAM-based tracking module terminates in the middle of the volume, we need to find turning points as new seeds to track segments along other directions.
+Naively, we can densely sample points from the segmentation result and run the plane-select module to find points that prefer other directions.
+However, such an approach can be inefficient as most points may prefer the original tracking direction, and ineffective due to false turning points caused by SAM errors.
+Instead, we design this module to sample turning points around the last point prompt position from the SAM-based tracking module.
+We first predict 2D SAM segmentation $seg_{p}$ along the other two directions ($p\in\{x,y,z\}\setminus\{p^{*}\}$) at the last point prompt position, which has a high probability of capturing the tangent cross-section of the swerving blood vessel.
+Then, we sample $K$ points from each of the 2D segmentation $seg_{p}$ with the Farthest Point Sampling (FPS) method as turning points added to the seed list.
+$seeds=\bigcup_{p\in\{x,y,z\}\setminus\{p^{*}\}}\mathrm{FPS}(seg_{p},K).$ (3)
+Algorithm 1 TriSAM Framework
+3D image volume $I$ and initial seed $s_{0}$, threshold $\tau$
 […]
@@
 […]
-     if $\operatorname{Valid}(Seeds)$ then
-         $SeedTree=\operatorname{LeafSplit}(SeedTree,{\textnormal{t}}_{i},Seeds)$triangleright Grow the seed tree
+     $p^{*}=\textsc{plane-select}(I,s)$
+     if not $\textsc{Visited}(s,p^{*})$ then
+         $seg=\textsc{SAM-Track}(I,s,p^{*})$
+         if $seg\neq\phi$ then
+              $seeds=\textsc{TurningPoint-Sample}(seg,p^{*})$
+[…]
+[…]
          end if
-end for
+     end if
+end while
 […]
+[…]
+4.3 Implementation Details
+Image Pre-processing.
+The large-scale datasets have many image artifacts and missing slices, which degrade the SAM-based tracking results.
+We utilize the temporal smoothing method along the z-axis to deflicker the images.
+Initial seed generation.
+Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background.
+To improve efficiency, we only keep the center of each connected component as the final seeds.
+SAM details.
+Unless stated otherwise, we use MobileSAM [36] instead of the standard SAM [15] in all conducted experiments to improve the inference speed.
+To clean up SAM results, holes are filled and small connected components are removed through binary morphological operations.
+Hyperparamter selection.
+As a zero-shot approach, TriSAM sets the hyperparameters to reasonable values instead of using ground truth data.
+In later ablation studies, we verify the robustness of hyperparameters within the reasonable range.
+For the global color thresholding for seed generation, $\eta$ is set to 98 percentile.
+For SAM-based tracking, the probability threshold $\tau$ is set to 0.8.
+Inference.
+For BvEM-Macaque, we directly predict the whole volume.
+For others, $k\times 1024\times 1024$ subvolumes are used since 1024$\times$1024 is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse.
+The subvolume results are later fused to form the final prediction.
+All experiments are conducted on an NVIDIA-A100 GPU.
 5 Experiments
 5.1 Experimental Settings
-Implementation Details.
-Since the volume is too large, non-overlapping subvolumes are used for segmentation and the results are fused to form the final prediction.
-Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background.
-Therefore, pixels darker than a threshold are considered as background while others are seed candidates.
-To improve efficiency, we only keep the center of each connected component as the final seeds.
-For BvEM-Macaque, we directly predict the whole volume.
-For others, $k\times 1024\times 1024$ subvolumes are used since it is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse.
-For preprocessing, we utilize deflickering method and temporal smoothing on the z-axis to remove the artifacts and missing slices.
-For postprocessing, holes are filled and small connected components are removed after the segmentation.
-Global color thresholding is used to generate the initial seeds and $\eta$ is set to 0.98 as shown Fig.
-8 (left).
-Given the significantly greater complexity of human blood vessels, we employ the manually generated seeds for the challenging BvEM-Human volume.
-For SAM-based tracking, the probability threshold $\tau$ is set to 0.8.
-according to the experiment shown in Fig.
-8 (right).
-Unless explicitly stated otherwise, we use MobileSAM [35] instead of the standard SAM [15] in all conducted experiments to improve the inference speed.
-All experiments are conducted on an NVIDIA-A100 GPU.
 Evaluation Metrics. We use the Precision, Recall, and Accuracy metric defined in [46], where $Accuracy=\frac{TP}{TP+FP+FN}$.
@@
 The compared zero-shot methods are global color thresholding and SAM+IoU tracking.
-For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10x512x512).
+For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10$\times$512$\times$512).
 Then we label all voxels that are 3 standard derivations above mean as positive.
@@
 The final instance segmentation results on BvEM-Macaque are shown in Fig.
-6.
+5.
 Color thresholding segments bright pixels, inadvertently capturing nuclei cells while overlooking darker pixels corresponding to blood vessels.
@@
 TriSAM prediction emerges as the most accurate method, affirming its effectiveness.
+Figure 6: Results with oracle 2D segmentation.
+Figure 7: Results for different plane selection strategies without recursive seed sampling.
+Table 3: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
+Method Pre. (%) Rec. (%) Acc. (%) Speed (sec)
+Single-plane (z) 75.28 48.48 41.82 324
+Single-plane (y) 79.37 58.47 50.75 307
+Single-plane (x) 69.73 12.14 11.53 345
+Single-plane (fusion) 71.78 74.15 57.41 976
+Tri-plane 78.41 74.97 62.14 335
 5.3 Ablation Studies
 We conducted a comprehensive series of ablation studies exclusively using the BvEM-Macaque dataset due to the computation constraints.
-Effect of the Tri-plane Selection.
+Oracle analysis.
+We conduct an oracle analysis to showcase the advantage of the proposed TriSAM framework for its capability to deal with complicated vascular geometry.
+For the SAM-based tracking module in TriSAM, we plug in an oracle 2D segmentation method which returns the connected component of the ground truth mask containing the prompt.
+As shown in Fig.
+7b, the TriSAM framework can perfectly segment the whole vasculature with the oracle 2D segmentation.
+In comparison (Fig.
+7a), existing SAM-based video object tracking methods can only track along one axis, missing bifurcated branches.
+Tri-plane selection module.
 We first compare our method with the single-plane methods to evaluate the effectiveness of the tri-plane.
@@
 Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.
-Table 3: Ablation studies on different plane selection strategies. The proposed tri-plane approach achieves the best overall accuracy with comparable speed.
-Method Pre. (%) Rec. (%) Acc. (%) Speed (sec)
-Single-plane (z) 75.28 48.48 41.82 324
-Single-plane (y) 79.37 58.47 50.75 307
-Single-plane (x) 69.73 12.14 11.53 345
-Single-plane (fusion) 71.78 74.15 57.41 976
-Tri-plane 78.41 74.97 62.14 335
-Figure 7: Segmentation results with different plane selection strategies.
-Figure 8: Ablation study results on hyperparameters $\eta$ and $\tau$.
 Figure 8: Histogram of the tri-plane selection.
@@
 We visualize the segmentation results with one initial seed on BvEM-Macaque in Fig.
-8.
+7.
 The origin point is the seed location.
@@
 However, it still falls short of the proposed method’s performance, as the latter takes into account potential turning points to exploit the 3D structure and perform long-term tracking.
-Table 4: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
-Strategy Pre. (%) Rec. (%) Acc. (%) Speed (sec)
-Naive 79.37 58.47 50.75 307
-Dense sampling 86.94 24.84 23.95 2001($\uparrow$ 552%)
-Recursive sampling 78.41 74.97 62.14 335 ($\uparrow$ 9%)
-Effect of the Recursive Redirection Strategy.
+Table 4: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
+Backbone Pre (%) Rec (%) Acc (%) Speed Parameters
+SAM [15] 77.65 66.03 55.48 1535s 615M
+MobileSAM [36] 78.41 74.97 62.14 335s 9.66M
+SAM-based tracking module.
+We examine the effect of different SAM variants, i.e., MobileSAM, on the TriSAM performance.
+As shown in Tab.
+4, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed.
+Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.
+Turning point sampling module.
 To perform long-term tracking and fully leverage the 3D blood vessel structure, we introduced recursive seed sampling by considering potential turning points.
@@
 This could be attributed to the frequent axis changes potentially leading to the omission of certain parts of the blood vessel and causing splitting errors.
-Table 5: Ablation study results on the choice of SAM models. The MobileSAM model achieves better performance with faster inference speed and smaller model size.
-Backbone Pre (%) Rec (%) Acc (%) Speed Parameters
-SAM [15] 77.65 66.03 55.48 1535s 615M
-MobileSAM [35] 78.41 74.97 62.14 335s 9.66M
-Quality Analysis.
+Table 5: Ablation studies on different seed sampling strategies. The proposed recursive seed sampling approach achieves the best overall accuracy with comparable speed.
+Strategy Pre. (%) Rec. (%) Acc. (%) Speed (sec)
+Naive 79.37 58.47 50.75 307
+Dense sampling 86.94 24.84 23.95 2001($\uparrow$ 552%)
+Recursive sampling 78.41 74.97 62.14 335 ($\uparrow$ 9%)
+Tri-plane dynamics.
 We delve deeper into the plane dynamics in Fig.
-11 and 11.
+10 and 10.
 The majority of the selected plane predominantly tracks along the y-axis.
@@
 In Fig.
-11, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses.
+10, we explore the size dynamics using various methods, where the mean derivation is shown in parentheses.
 The size variation observed with our proposed method is relatively smaller compared to tracking along the y-axis and significantly less than when tracking along the z-axis and x-axis.
-Other Ablations.
-To further improve the efficiency, we use MobileSAM instead of the original SAM in practice.
-As shown in Table 5, the inference time of MobileSAM is 22% of the original SAM which confirms that MobileSAM significantly improves the inference speed.
-Moreover, the performance of MobileSAM is even better than the original SAM, possibly because the distilled small model is less prone to overfitting to the original natural image domain.Fig.
-8 illustrates the performance with varying hyperparameters $\eta$ and $\tau$.
-Figure 12: The failure cases of the proposed method. Three prediction and ground truth pairs are shown.
+Figure 11: Ablation study results on hyperparameters $\eta$ and $\tau$.
+Hyperparameters.
+We compute the accuracy of TriSAM with varying hyperparameters for seed generation ($\eta$) and SAM-based tracking module ($\tau$).
+As shown in Fig.
+11, TriSAM’s performance is robust within the reasonable range of hyperparameters.
 5.4 Failure Cases
-Despite our method’s significant progress, it still encounters challenges in some scenarios, as illustrated in Figure 12.
-The first issue is the seed problem, where the presence of a false positive seed can lead to erroneous predictions.
-Secondly, SAM could also fail to segment the mask.
-Lastly, tracking failure tends to occur at blood vessel conjunction points.
-These instances highlight the complexity of tracking in such intricate vascular networks and suggest the need for further refinements in our approach to address these specific failure cases effectively.
+The main source of error in the proposed TriSAM framework is the SAM-based tracking module, where SAM can not produce accurate 2D mask given the prompt.
+For the false negative case (Fig.
+12a), SAM predicts an empty mask when the shape and appearance of the blood vessels are complex (e.g., conjunction point) for all three axes.
+For the false positive case (Fig.
+12b), SAM falsely segment the neighboring region with the blood vessel due to the similar image appearance.
 6 Conclusion
@@
 By addressing a significant gap in the field of neuroimaging, we have laid the foundation for advancing the understanding of cerebral vasculature at the microscale and its intricate relationship with neural function.
+Figure 12: Failure cases of the SAM-based tracking module.